### v1

# Diffusion in Zero-Shot Learning for Environmental Audio

Abstract
Zero-shot learning enables models to generalize to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes. While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies. Generative methods, which have demonstrated success in computer vision, are notably absent from environmental audio zero-shot learning, where classification-based approaches dominate.
To address this gap, this work investigates generative methods for zero-shot learning in environmental audio. Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN). Additionally, a novel diffusion model conditioned on class auxiliary data is introduced. The diffusion model generates synthetic data for unseen classes, which is combined with seen-class data to train a classifier.
Experiments are conducted on two environmental audio datasets, ESC-50 and FSC22. Results show that the diffusion model significantly outperforms all baseline methods, achieving more than 25% higher accuracy on the ESC-50 test partition.
This work establishes the diffusion model as a promising generative approach for zero-shot learning and introduces the first benchmark of generative methods for environmental audio zero-shot learning, providing a foundation for future research in the field.
Code is provided at https://github.com/ysims/ZeroDiffusion for the novel ZeroDiffusion method.
Index Terms:
Environmental Audio, Zero-Shot Learning, Generative Networks
I Introduction
Machine learning has revolutionised how society solves complex problems [1, 2, 3]. Various publications consider the future impact of machine learning, such as AI cities [4], humanoid robotics [5] and medicine [6, 7, 8]. To solve these future problems, machines and AI need to be adaptive [9, 10]. One way machines can be adaptive is to classify new items without labelled training data [11, 12]. This problem is called ‘zero-shot learning’.
Zero-shot learning is a category of supervised machine learning that aims to create a model to classify classes it has not seen during training. That is, the set of classes used in training and testing are disjoint. Zero-shot learning is valuable in real-world applications, such as remote sensing [13], image classification [14, 15], image segmentation [16, 17], environmental audio [18, 19] and security [20], where there may not exist data for every class.
An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem. Suppose someone has never seen a snake, cat or bird but knows key attributes these animals have. When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings. The person performs zero-shot learning using features of the target class and learning what those features mean through knowledge of other animals.
It is important not to confuse zero-shot learning with similar zero-data problems. In zero-shot learning, the method knows about the existence of all classes, i.e. the set of classes is closed. If it was open, i.e. there are unknown classes, the problem is called open-set learning [21, 22, 23]. Domain adaptation [24, 25] is another problem with a lack of data; however, there is no clear ‘seen’ and ‘unseen’ class distinction. An example of domain adaptation is training in simulation and transferring to the real world without fine-tuning [26, 27, 28]. Finally, zero-shot prompting is a newer area that uses large language models to perform new tasks [29, 30, 31]. This differs from zero-shot learning as there is no seen and unseen class distinction, and it is hard to quantify how much overlap the task has to scenarios the model was trained on.
Close to these zero-data problems, particularly zero-shot learning, is few-shot learning. Rather than zero data for select classes, few-shot learning has a small number of samples (e.g. three to five) for particular target classes. A comprehensive review of few-shot learning is provided by Song et al. [32].
In the literature, zero-shot learning methods are commonly tested on computer vision classification tasks [12, 33, 34, 14, 35, 36, 15, 37, 38, 39]. Traditional methods for zero-shot learning in computer vision relied on classification-based models with engineered attributes [12], but more recent methods use generative models [34, 15, 37] such as generative adversarial networks [40] and variational autoencoders [41]. These models have been shown to improve performance on unseen classes and can help mitigate problems such as the hubness problem and the domain shift problem [42]. The hubness problem occurs when data points are concentrated around a minority of classes, leading to poor performance when tested on unseen classes [43]. The domain shift problem arises when the model performs poorly because the training and test sets have different distributions [44].
This work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision. It takes inspiration from the successful usage of generative methods in computer vision [14, 45, 34, 37], and from a newer generative method, diffusion, which has seen success in image generation tasks [46, 47]. There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
Recent use cases for environmental audio classification include smart cities [48], detection of traffic noise pollution [49], and hearing aid analysis [50]. These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
There is a lack of datasets that can be used to benchmark environmental audio zero-shot learning, as datasets are often not released in their raw audio form or lack classes. AudioSet [51] is a valuable large dataset, but instead of audio clips, it contains embeddings. Embeddings are high dimensional vectors representing the data. These embeddings come from a network pre-trained on AudioSet. This pre-training makes the dataset unsuitable for zero-shot learning, as the unseen classes may be present. Other datasets, such as UrbanSound8k [52], have too few classes for rigorous zero-shot learning. Partitioning seen and unseen classes will result in a seen set with little diversity and may not cover a large enough semantic space for the model to generalise to unseen classes, risking overfitting. Additionally, the unseen classes will be very few, and so the statistical significance of the results is poor. A reasonable minimum may be five unseen classes, and double that amount of seen classes, for a total of 15 classes in the dataset.
ESC-50 [53] is an open dataset with fifty classes, often used in environmental audio zero-shot learning literature. FSC22 [54] is another open dataset and contains twenty-seven classes. While this is almost half the size of ESC-50, it is still a substantial size to perform zero-shot learning on. We are not aware of the existence of any zero-shot learning literature using this dataset.
The present study uses ESC-50 and FSC22 to compare three generative methods. These methods use the following architectures: generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion networks. A bilinear compatibility network [55] is also compared as a baseline. The method ZeroDiffusion is introduced as a new method for zero-shot learning using diffusion.
The GAN method used is LisGAN [15]; however, the component where high-confidence unseen samples are injected back into the training is not included, to prevent violating the zero-shot learning problem formation where unseen class data should not be present during training. The method uses a conditional Wasserstein GAN [56]. Zero-shot learning methods often aim to learn the relationship between feature and class embeddings. To achieve this, the GAN is conditioned on class embeddings. It is then used to generate synthetic samples for unseen classes. Vanilla conditional GAN methods were experimented with in the presented study, but they suffered extreme instability and low performance. The adjusted LisGAN method performs worse than the bilinear compatibility method but is better than a random model.
The cross-aligned and distribution-aligned VAE (CADA-VAE) [14] has been previously used in computer vision with success. The method trains two variational autoencoders, one for auxiliary data and one for feature data, to learn a shared latent space. The results were comparable to the bilinear compatibility method.
Finally, the most successful model is our new diffusion model. The experiments with diffusion show a clear improvement upon any previous method in environmental audio zero-shot learning. While there is existing zero-shot learning literature outside of environmental audio on diffusion models [57, 58], no method trains a network from scratch. The methods use the pre-trained Stable Diffusion model [47]. While this can provide practical usage, it is hard to draw theoretical conclusions on the zero-shot performance of this method. The overlap of the unseen classes and training classes is unknown. This work shows that diffusion, with a transparent training process and a valid seen and unseen class distinction, outperforms other techniques in zero-shot learning.
The contributions of the presented paper are:
  • The first diffusion method for zero-shot learning that does not use a pre-trained model.
  • A new state-of-the-art for zero-shot learning with the ESC-50 dataset using the diffusion method.
  • The first benchmark on the FSC22 dataset, with the diffusion method outperforming other methods.
  • The first comparison of GANs, VAEs and diffusion-based methods for environmental audio zero-shot learning with ESC-50 and FSC22.
II Related Work
This section first covers terminology in zero-shot learning to better understand the literature. Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision. Finally, existing works in environmental audio zero-shot learning are considered.
II-A Terminology
In this section, zero-shot learning terminology is presented in an effort to contextualise the literature. First, embeddings for both class and feature data are considered, then the concept of auxiliary data and finally, a high-level definition of zero-shot learning.
II-A1 Embeddings
Embeddings are a method of converting raw data, such as images or audio, into a high-level form that machine learning algorithms can use. The form embeddings take is a high-dimensional vector of real numbers. The methods in this work use deep learning to obtain these vector representations, using both supervised and unsupervised methods. An example of what an embedding space might look like is shown in Figure 1. Lions and cats are both felids and therefore semantically similar and closer to each other in the vector space. ‘Dog’ is semantically closer to the felids than ‘car’ as they are types of animals. In this way, the structure of the space reveals semantic relationships.
Figure 1: An embedding space. Semantically similar data is closer, such as ‘lion’ and ‘cat’.
Supervised methods to create embeddings typically involve classical supervised learning on a dataset and the output of the second last layer of the artificial neural network becomes the embedded data vector. Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set. Typically the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset could be used as long as there is no overlap with the unseen class set.
Unsupervised methods involve learning a mapping to a space where the semantic distance of data is represented in the space. The most common example of this embedding method is word embeddings, such as GloVe [59] and Word2Vec [60]. These methods learn the semantic similarity of words using their proximity in texts. To a human, embeddings may seem to contain little meaning, but these vectors are information-rich. For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [60].
II-A2 Auxiliary Data
Zero-shot learning involves learning to classify known classes without samples for those classes. If the classes are unknown, the problem is called open-set learning. These known classes contain information even without samples. This information is called ‘auxiliary data’. The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe. Another popular form is human-engineered attributes, where a vector is created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’. The Animals with Attributes [12] and Animals with Attributes 2 [61] datasets are examples of datasets that contain human-engineered attribute vectors, and were created for zero-shot learning research.
II-A3 Zero-Shot Learning
Zero-shot learning involves training on seen classes and testing on unseen classes. Generalised zero-shot learning extends this problem to include both seen and unseen classes in testing. Zero-shot learning is one of many zero-data problems, as described in the introduction.
Feature data is defined as the target data the user wants to classify. In computer vision, the feature data is images. In environmental audio, it is audio clips. Feature data for unseen classes should not be present anywhere before the testing phase, including during pre-training.
The embeddings for the feature data and classes represent semantic information. Zero-shot learning aims to train a model that generalises its understanding of semantics to go beyond class-specific knowledge.
II-B Generative Methods in Zero-Shot Learning
While there is no literature known to the authors in environmental audio that considers generative networks, there are many examples in the domain of computer vision. These largely involve networks combining GANs and autoencoders.
In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al. [33] and Mukherjee et al. [62]. These used two traditional autoencoders for each modality and learnt a shared latent space between these two autoencoders. Conditional variational autoencoders were introduced in zero-shot learning the following year [34]. A new benchmark was found when CADA-VAE [14] was proposed which took advantage of variational autoencoders and created a shared latent space through cross-alignment and distribution-alignment.
One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [36]. In the same year, another GAN method was created that used a visual pivot regulariser to guide training [63]. It also used Wikipedia text descriptions converted to word embeddings using term frequency-inverse document frequency (TF-IDF). The method was compared with other zero-shot learning methods including a basic GAN model. The basic GAN model performs much worse than any established zero-shot learning method, while Zhu et al.’s method [63] shows an improvement over the state-of-the-art.
LisGAN [15] improved on previous GAN methods by using a Wasserstein GAN conditioned with auxiliary data. They use ‘soul samples’, samples representing the ‘ideal’ sample from a class. Soul samples for seen classes are the average sample over all the seen samples for that class, while soul samples for unseen classes are constructed using the auxiliary data for the class. The GAN is trained to generate data close to a soul sample. The authors perform a step where unseen data is tested and any samples with high confidence are injected into training. This violates the requirement in zero-shot learning where unseen data should not be seen during training, even if it is unlabelled. GDAN [37] proposes a GAN with a regressor and cyclic-consistent loss. The results are marginally better than the conditional VAE method from Mishra et al. [34], indicating that GAN methods still fall behind VAE methods for zero-shot learning, given that at this point better VAE methods had been introduced.
Task Aligned Generative Meta-learning for Zero-shot Learning [38] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes. The backbone of the technique is a meta conditional GAN [64]. SAGAN [39] involves training a GAN, followed by training a classifier to align the visual and semantic space. This method considers what to do with the trained GAN rather than improving the GAN training method. In the same year, ZeroNAS [65] is proposed which performs neural architecture search on GAN zero-shot learning methods to find the optimal architecture for a zero-shot learning problem. The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of the expert curation of these networks.
While there is much literature on how GANs may be used for zero-shot learning and how general improvements to GAN training can be used, they are still notoriously hard to train [63, 65]. This does pose the question of if GANs are too unstable and not reliable enough for zero-shot learning problems. In practical situations, it is hard to estimate model performance on unseen classes. Therefore, the method must be reliable.
Chou et al. [35] propose a generative method without VAEs or GANs. It uses a beta distribution to generate virtual classes for more robust zero-shot learning. Despite their efforts, GANs and VAEs still dominate the landscape of zero-shot learning.
In 2023, two publications emerged using diffusion in zero-shot learning [57, 58]. These both exist within the image classification domain and use the pre-trained Stable Diffusion model [47]. These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning.
II-C Environmental Audio Zero-Shot Learning
Limited research exists for environmental audio zero-shot learning. One of the earliest works proposes a Twin neural network [66], tested on a small dataset containing ten classes. Xie and Viranen [18] used a more substantial dataset of fifty classes. In addition to a bilinear compatibility model [67] that created a baseline for the ESC-50 dataset, they proposed two partition strategies for five-fold cross-validation. These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
Further work in the field [68, 69] brought new state-of-the-art results and critical considerations when training zero-shot networks. Xie and Viranen [68] highlight the importance of the unseen and seen class distinction, where they identify that VGGish pre-trained on AudioSet is unsuitable for zero-shot learning on ESC-50 due to the presence of unseen classes in AudioSet.
AudioSet contains embeddings rather than audio clips. Embeddings are a critical aspect of zero-shot learning methods, and some research works explore improving these embeddings [19]. Furthermore, it is vital to maintain the distinction between ‘seen’ and ‘unseen’ classes. Therefore, training from raw data is essential. AudioSet provides the location of the YouTube videos the clips come from, but it is against the Terms of Service to download the raw data. AudioSet is, therefore, unsuitable for zero-shot learning research.
Some research focuses on improvements through embeddings. While not often the main focus of methods, they are the foundations of the zero-shot method. The remaining pipeline is limited to the information that the embeddings can capture. The methods in this paper use work by Sims et al. [19] to improve the performance through embeddings by adding synonyms.
III Method
This section will begin with a technical overview of zero-shot learning to complement the high-level background given in the previous section. Following this background information, the novel zero-shot learning diffusion method (ZeroDiffusion) is presented. A visual description of the method is shown in Figure 2.
III-A Zero-Shot Learning
The zero-shot learning problem is defined as the following. There exists a set $X$ of seen data and a set $Y$ of unseen data such that $X\cap Y=\varnothing$. The zero-shot learning method aims to learn a function
$f(z,C)=c$ (1)
where $z\in X\cup Y$ is a data sample; $C$ is the set of all classes; and $c\in C$ is the class that the sample $z$ comes from. After training a model to learn $f$ using all samples in $X$, the model is tested on all samples in $Y$.
The literature frequently divides zero-shot learning methods into three components. One component is the auxiliary data function $\sigma(c)$ that takes a class $c\in C$. The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$. $\sigma$ changes a class label into high-level information.
A second component is the feature embedding function $\tau(z)$, where $z\in X\cup Y$ is any feature data in the dataset. The feature data in the case of environment audio is audio clips. $\tau$ will convert the feature data into high-level information, similarly to $\sigma$. The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer. In computer vision, a popular method is to use ResNet, pre-trained on ImageNet.
The third component ties the previous two together to retrieve a class prediction. This component learns a function
$\xi(\tau(z),\sigma(C))=c$ (2)
that takes the high-level information output of the other two components and outputs the corresponding class $c$ for a sample $z$. This component of zero-shot learning is explored heavily across the literature.
III-B Embeddings
An embedding is an encoded form of data that is easy for a machine learning model to process. They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
In the context of zero-shot learning, embeddings are beneficial when they contain semantic information about the data. Learning the relationship between the feature and class semantic embeddings is crucial to performing zero-shot learning.
The $\sigma$ function is the Word2Vec [60] pre-trained model. This model takes class labels and outputs an embedding vector for that word. The space that Word2Vec has learned is information-rich, with semantically similar words existing closer together in the space than semantically distant words. Word2Vec gives an embedding in $\mathbb{R}^{300}$.
We use work by Sims et al. [19] to enrich the word embeddings with synonyms. The average word embedding of the class label and synonyms are used as auxiliary information. For ESC-50, the synonyms from the original paper were used. FSC22 has not been used in previous literature on synonyms, and so we human-engineer the synonyms, where overlapping classes use the same synonym list from ESC-50.
To learn $\tau$ and obtain audio embeddings, we train a modified YAMNet model, shown to perform better than other methods [19], on the seen class data. As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
Thus both $\sigma$ and $\tau$ have been defined, where the output of $\sigma(c)$ is a class embedding of dimension $300$ for the class $c$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w\in X\cup Y$.
III-C ZeroDiffusion
Figure 2: ZeroDiffusion involves training three models. The first is the audio embedding model, used to retrieve audio embeddings. Seen classes are used to train this model. Next, the diffusion model is trained using audio embeddings from the first model and word embeddings. This process is done using seen classes. Finally, a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
Diffusion involves training a model to reconstruct data from noise. During training, the input to the model is a data sample with noise added, and the expected output is the original data sample. The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise. This process gradually trains a model to generate data from noise.
Conditioning a diffusion model allows the user to control the output. Image generation tasks condition the diffusion model on textual data descriptions. This conditioning allows the user to give a trained diffusion model a description of any image they want to generate [70]. If a user requests output of ‘a cat holding a basketball’, the model should be able to recreate a picture of a cat holding a basketball from noise.
A diffusion model conditioned on auxiliary data is trained with the input to the model being paired embeddings $\sigma(c)$ and $\tau(w)$ where $c$ is the corresponding class for the audio sample $w\in X$. The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(c)$.
The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with mean 0 and variance 0.1 to create a gradually noisy embedding vector
$\tilde{\tau}(w)(p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
where $p$ is the epoch percentage.
The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(c)$. The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428. Due to the concise and information-rich nature of embeddings, the diffusion architecture is small and uses fully connected layers. The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size. This is followed by batch normalisation for training stabilisation and then a Leaky ReLU activation to add a non-linearity component. Following Leaky ReLU, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space. The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
The loss is the mean squared error between the embedding the model is trying to recreate and the output of the model. During training, gradients are clipped to a maximum norm of 1 to stabilise training. An Adam optimiser is used.
Our new diffusion model is trained on seen class data. The aim is for the model to generalise to generating unseen data. Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise. After creating an unseen dataset in this manner, a bilinear compatibility network [67, 18] is trained with both the true seen data and synthetic unseen data. The bilinear network uses a cross-entropy loss function and an Adam optimiser.
IV Experiments
Experiments are performed on the ESC-50 and FSC22 datasets. The state-of-the-art in zero-shot learning environmental audio, the compatibility bilinear method [67, 18] is compared with existing methods from computer vision [14, 15], and our new method ZeroDiffusion.
ESC-50 is split into five partitions of ten classes. One partition is held as the test set. The remaining four partitions are used for four-fold cross-validation, where three of the partitions are used in training and the remaining one is used for validation. The validation set is used to tune the model and its hyperparameters before using the test set. This design is modified from the five-fold cross validation used in previous work [68, 19] to include a held-out test set. Using a four-fold strategy for hyperparameter tuning allows for more generalised hyperparameter tuning and a better understanding of the performance of the model. The final experiments on the held-out test set give a good indication of how the model performs on a completely unseen set of classes and data. ESC-50 is a suitable dataset for this strategy as it is composed of many classes, allowing for each partition to have ten classes. The dataset is correctly labelled and clean, making it ideal for testing.
FSC22 is split into three sets, a test set with 7 classes, a validation set of 7 classes, and a training set with the remaining 13 classes. The validation set is used for hyperparameter tuning, and then the method is evaluated using the test set. This dataset has almost half the number of classes compared to ESC-50, reducing the amount of flexibility in forming a strategy. Additionally, the number of classes is uneven and has few divisors. If it were partitioned into nine lots of three, the number of classes in each set would be extremely small for zero-shot learning while also creating a lot of folds to manage. If three lots of nine were used, the training set would be very small compared to the usual split in zero-shot learning, where there are more training classes than test classes. Therefore, the validation and test sets are chosen to be the same size for consistency, and train with a larger number, being almost the size of the test and validation sets combined. Therefore, $k$-fold validation is not performed for this dataset. The dataset does overlap with the classes in ESC-50, but there are new classes not present in ESC-50. Some classes have questionable labels, including the class ‘silence’, which appears to contain a variety of soundscapes including birds chirping and rain falling, which are classes elsewhere in the dataset. Because the dataset is not entirely well-labelled, this is considered a harder dataset compared to ESC-50.
All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were run on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
Index Class Index Class Index Class
0 Dog 17 Pouring water 34 Can opening
1 Rooster 18 Toilet flush 35 Washing machine
2 Pig 19 Thunderstorm 36 Vacuum cleaner
3 Cow 20 Crying baby 37 Clock alarm
4 Frog 21 Sneezing 38 Clock tick
5 Cat 22 Clapping 39 Glass breaking
6 Hen 23 Breathing 40 Helicopter
7 Insects (flying) 24 Coughing 41 Chainsaw
8 Sheep 25 Footsteps 42 Siren
9 Crow 26 Laughing 43 Car horn
10 Rain 27 Brushing teeth 44 Engine
11 Sea waves 28 Snoring 45 Train
12 Crackling fire 29 Drinking, sipping 46 Church bells
13 Crickets 30 Door knock 47 Airplane
14 Chirping birds 31 Mouse click 48 Fireworks
15 Water drops 32 Keyboard typing 49 Hand saw
16 Wind 33 Door, wood creaks
TABLE I: Class indices and names in ESC-50. This follows the convention from the dataset.
Index Class Index Class Index Class
0 Fire 9 Axe 18 Footsteps
1 Rain 10 Chainsaw 19 Clapping
2 Thunderstorm 11 Generator 20 Insect
3 Water Drops 12 Handsaw 21 Frog
4 Wind 13 Firework 22 Bird Chirping
5 Silence 14 Gunshot 23 Wing Flapping
6 Tree Falling 15 Wood Chop 24 Lion
7 Helicopter 16 Whistling 25 Wolf Howl
8 Vehicle Engine 17 Speaking 26 Squirrel
TABLE II: Class indices and names in FSC22. While the same order is followed from the dataset, here the classes are zero-indexed rather than one-indexed.
Partition Classes
ESC-50 fold 0 2, 3, 27, 29, 31, 35, 38, 40, 46, 48
fold 1 13, 19, 21, 22, 26, 32, 36, 39, 42, 49
fold 2 4, 10, 14, 17, 23, 24, 30, 33, 41, 45
fold 3 1, 6, 7, 18, 20, 25, 28, 34, 44, 47
fold 4 0, 5, 8, 9, 11, 12, 15, 16, 37, 43
FSC22 train 0, 1, 2, 3, 4, 10, 11, 14, 16, 19, 20, 24, 25
val 6, 8, 9, 12, 13, 18, 22
test 5, 7, 15, 17, 21, 23, 26
TABLE III: Classes in each partition. For ESC-50, random partitions are used from previous work for consistency [18]. FSC22 has not been benchmarked for zero-shot learning before, so a new partition strategy is created with a random number generator.
A modified YAMNet model [19] is trained using only the seen classes in FSC22. A learning rate of $10^{-3}$ is used and early stopping occurred at epoch 259 with an accuracy of $0.6833$. This model is used to retrieve the audio embeddings. The word embedding model is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/. Each method uses the same embedding vectors for each respective class and data sample.
IV-A Bilinear Compatibility
The bilinear compatibility network has been used to great success in environmental audio [18, 68, 19]. It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50. Therefore, it is important to include it in this comparison.
The bilinear compatibility method has few hyperparameters and the parameters used by Xie and Virtanen [18] for ESC-50 were repeated in our work. FSC22 was tuned to train for 30 epochs.
IV-B CADA-VAE
Cross and distribution aligned variational autoencoder (CADA-VAE) [14] is a successful zero-shot learning method in computer vision, and for this reason it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets. Both datasets used the same hyperparameters, aside from the number of classifier epochs. Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy. It was found that the large amount of hyperparameters made this method difficult and time consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error. The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22. The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs.. The classifier has a learning rate of $10^{-3}$, a batch size of 32 and is trained for 23 epochs for ESC-50 and 20 epochs for FSC22. The loss of the VAE is mean absolute error and the loss of the classifier is negative log likelihood loss. Both use an Adam optimiser. The size of the latent space is 64.
IV-C LisGAN
Leveraging invariant side generative adversarial network (LisGAN) [15] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning. While there are some methodological concerns, these were modified to adhere to the zero-shot learning problem formulation and the core idea of the method remained. A conditional Wasserstein GAN is the main GAN method found in zero-shot learning. However, despite efforts to improve the method, it is still notoriously difficult to train [71, 65]. Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs. LisGAN showed to be significantly more stable and competitive amongst other zero-shot learning methods.
Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune. These mostly were modified to suit the smaller ESC-50 and FSC22 datasets, as the datasets in computer vision are significantly larger. Both datasets used the same hyperparameters.
LisGAN is used without retraining on unseen classes. The GAN is trained with a batch size of 64, 128 units in the hidden layer of the discriminator and generator, 50 epochs and a learning rate of $10^{-4}$. The classifier is trained with a learning rate of $10^{-2}$ and 50 epochs. The generator, discriminator and classifier use Adam optimisers. The classifier uses a negative log likelihood loss.
ESC-50 FSC22
Val Fold 0 Val Fold 1 Val Fold 2 Val Fold 3 Test Val Test
Bilinear $0.2460\pm 0.0074$ $0.5003\pm 0.0074$ $0.3345\pm 0.0105$ $0.2868\pm 0.0076$ $0.2257\pm 0.0105$ $0.3526\pm 0.0069$ $0.3343\pm 0.0169$
LisGAN $0.2605\pm 0.0227$ $0.4039\pm 0.0220$ $0.2582\pm 0.0187$ $0.2150\pm 0.0195$ $0.2448\pm 0.0165$ $0.299\pm 0.0219$ $0.2142\pm 0.0416$
CADA-VAE $\textbf{0.2915}\pm 0.0180$ $0.3980\pm 0.0181$ $0.2500\pm 0.0251$ $0.3013\pm 0.0251$ $0.2095\pm 0.0337$ $0.3179\pm 0.0137$ $0.3290\pm 0.0339$
ZeroDiffusion $0.2846\pm 0.0111$ $\textbf{0.5279}\pm 0.0283$ $\textbf{0.3593}\pm 0.0189$ $\textbf{0.3342}\pm 0.0224$ $\textbf{0.3091}\pm 0.0278$ $\textbf{0.3975}\pm 0.0132$ $\textbf{0.3467}\pm 0.0200$
TABLE IV: Average accuracy and standard deviation over ten randomly seeded runs for the traditional bilinear compatibility method and the three generative methods for ESC-50 and FSC22. Our method, ZeroDiffusion, outperforms the other methods in accuracy across the partitions of both datasets. The bilinear compatibility method has less variance than the three generative-based methods.
IV-D ZeroDiffusion
ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in this paper. These experiments aim to determine the performance of this new method in the context of environmental audio. The existing success of diffusion in other areas of machine learning research motivate the creation and experimentation of this method.
The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 2000 epochs. Fifty samples are generated from the diffusion model per unseen class. The bilinear compatibility network is trained with a learning rate of $10^{-4}$, batch size of 52, and 20 epochs. These hyperparameters were chosen using manual tuning.
V Results
The results from experiments with the four methods - bilinear compatibility, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50 and FSC22 are presented in Table IV.
ZeroDiffusion increases the accuracy on the ESC-50 and FSC22 test partition by $36.95\%$ and $3.71\%$, respectively, from the baseline bilinear method. With the exception of the first validation fold in ESC-50, ZeroDiffusion has a higher accuracy than the other three zero-shot learning methods. This shows that ZeroDiffusion is a new state-of-the-art method for both ESC-50 and FSC22.
Notably, ZeroDiffusion incorporates the bilinear model in its methodology. The difference between the methods is that ZeroDiffusion trains a diffusion model to generate synthetic unseen data and adds these samples to the training of the bilinear method. Because the diffusion model is the only addition, it is reasonable to conclude that the diffusion model is responsible for the improved accuracy.
CADA-VAE has positive performance but is inconsistent across partition strategies. LisGAN has similar inconsistency issues but overall appears to have worse performance across partitions and datasets. This inconsistency continues when considering the standard deviation for each experiment. The generative methods have high variance compared to the baseline method. In the case of ZeroDiffusion, the higher average accuracy and variance are tradeoffs to consider.
The accuracy improvement of ZeroDiffusion supports the use of diffusion models in the zero-shot learning field. Adjacent works with Stable Diffusion [57, 58] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning.
Confusion matrices give more insight into model performance through the analysis of class predictions. The confusion matrices for one run of ESC-50 fold 0 and 1 are shown in Tables V and VI, respectively.
In fold 0 of the ESC-50 dataset, the helicopter and mouse click classes have the highest accuracy at 80% and 62.5%, respectively. The helicopter class is predicted at least 70% of the time for four different classes, resulting in a hub around the helicopter class. A hub is a class where predictions across classes are clustered, i.e. the model is much more likely to predict the hubs than other classes. In zero-shot learning, the hubness problem has been identified as a major issue preventing classes from being predicted, resulting in a low accuracy. In fold 0, smaller hubs exist around drinking, brushing teeth and mouse clicking. Other classes are not predicted often. The difference between the hub classes and other classes does not appear to have a clear intuitive explanation. This information could lead to avenues for improvement.
Fold 1 of ESC-50 has a significantly higher accuracy than fold 0. The highest class accuracy is siren at 80%, however all but the clapping class have at least 25% accuracy. Clapping is predicted a total of three times across all clips in this fold, and all three are misclassified hand saw clips. Other classes are predicted more consistently, indicating that the hubness problem isn’t as prevalent in the training of this partition of the dataset.
The confusion matrix in fold 1 is sparse, where classes are only misclassified as a small number of other classes. This phenomenon is also found in traditional supervised learning, where the target class is misclassified as classes with similar features. In zero-shot learning, it is the machine learning-based semantic embeddings that are being compared, and so intuition from traditional supervised learning may not apply.
Some misclassifications appear intuitive, such as the case where the sound of a washing machine is incorrectly labeled as a helicopter. Both classes are loud and mechanical, motivating a potential misunderstanding from the model. Fireworks are misclassified as mouse clicking in fold 0, and the reason may not initially be intuitive, but both sounds are comprised of short, sharp noises. There is a difference in loudness, but the audio data is normalised before input to the network and so this difference may be reduced. While there are some intuitive misclassifications, there are also non-intuitive misclassifications. Pig is incorrectly labeled as brushing teeth, and door knocking is mislabeled as rain.
To gain a deeper understanding, dimensionality reduction with t-distributed stochastic neighbour embedding (t-SNE) [72] on the audio embeddings produces Figure 3 for fold 0 and Figure 4 for fold 1. In fold 0, pig (class 2, dark purple) is spread across a wide area, including overlapping with brushing teeth (class 27, dark blue). This relationship gives insight into the non-intuitive misclassification of pig.
Fold 1 has a significantly higher accuracy than fold 0, as shown in Table IV. The t-SNE graph for fold 1 shows more distinct clusters than fold 0, indicating that this clustering property of the audio embeddings is important in correctly classifying unseen classes. This is reinforced when considering the clapping class, which both is the poorest performing class in fold 1 and has the greatest spread across the t-SNE graph.
The t-SNE graphs reveal insights into the hubness problem, and asymmetry in the incorrect predictions further highlights the problem. Pig is mislabeled as brushing teeth more than 50% of the time, while brushing teeth is never mislabeled as pig. This shows that certain classes are more likely to be predicted than others. The classes in fold 0 that are frequently predicted as the helicopter class are closer together in the t-SNE graph, showing their similarity in the embedding space. This similarity is a potential cause of the hubness problem.
Class index 38 29 35 3 40 2 27 46 31 48
38 1 9 0 3 0 3 9 0 14 1
29 1 18 0 0 1 0 18 0 1 1
35 1 0 6 0 30 1 1 1 0 0
3 0 0 5 0 29 5 1 0 0 0
40 0 0 6 0 32 0 0 1 0 1
2 0 6 3 0 3 5 21 0 0 2
27 2 34 0 1 0 0 1 0 2 0
46 0 0 2 0 28 2 0 8 0 0
31 1 5 0 0 2 0 6 0 25 1
48 4 6 0 0 2 0 8 1 15 4
TABLE V: Confusion Matrix for fold 0 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. Many predictions are clustered around class 40, with smaller hubs on classes 29, 27 and 31.
Figure 3: ESC-50 fold 0 audio embeddings with t-SNE dimensionality reduction. Class 40 (helicopter), 46 (church bells), 35 (washing machine) and 3 (cow) are all concentrated to the right side of the scatterplot. The confusion matrix in Table V shows that all of these classes are classified as helicopter a majority of the time.
Class index 39 36 42 13 32 22 19 49 26 21
39 10 0 0 0 9 0 2 3 0 16
36 0 24 3 4 0 0 6 1 0 2
42 0 0 32 3 0 0 0 2 3 0
13 0 4 7 14 1 0 5 0 1 8
32 2 0 0 0 23 0 2 13 0 0
22 14 1 0 1 3 0 14 6 1 0
19 0 5 10 0 0 0 22 0 0 3
49 1 3 0 1 0 3 0 11 8 13
26 4 0 2 4 0 0 0 4 15 11
21 2 0 0 1 1 0 0 0 8 28
TABLE VI: Confusion matrix for fold 1 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. There are small hubs on classes 42, 19, 26 and 21 and all classes are at least partially correctly identified, except for class 22 which is never correctly classified.
Figure 4: ESC-50 fold 1 audio embeddings with t-SNE dimensionality reduction. Fold 1 exhibits more defined class clusters than fold 0. Class 22 (clapping, dark blue) has a greater spread, and is the poorest classified class as shown in Table VI.
t-SNE graphs give a deeper insight into issues surrounding zero-shot learning, and can indicate avenues for improvement. The relationship between the clustering properties of the audio embeddings and model accuracy show that embeddings have an effect on the performance of a zero-shot learning model. More defined clustering in the audio embeddings produces higher accuracy, and hubs in the confusion matrix are reflected in the t-SNE diagram. The confusion matrices highlight the abilities of the model beyond an accuracy value. The sparsity of the matrices demonstrate that the model recognises the similarity of clips from the same class, and when it correctly maps these to the correct class the performance is very high.
VI Discussion
In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples. Therefore, consistency across training runs is important. While ZeroDiffusion has a superior accuracy-based performance, the bilinear method has superior consistency. This is an important consideration when choosing an appropriate method for zero-shot learning in practical applications.
The hubness problem is an issue when considering practical usage because of the poor accuracy on specific individual classes that are overshadowed by the hubs. Many of these specific classes are predicted at a rate worse than a random model. Hubness appears to correlate to overall model performance, as seen when comparing the confusion matrices and overall accuracy of two partitions of the ESC-50 dataset.
One key avenue in reducing the hubness problem is first understanding why the different partitions used in this work had varying levels of hubness. Partitions can be crafted for this purpose to identify the effect of different criteria on hubness. One question is whether the unseen classes, seen classes, or both, influence the performance. If particular characteristics in the seen class set influence the hubness problem, then the seen class set can be supplemented with classes exhibiting these characteristics. A more likely scenario is that both influence the performance, and the seen classes used in training may need to be curated for the characteristics of the specific unseen set in the given problem.
Zero-shot learning lacks theoretical foundations, with methods developing through empirical results. There is little understanding of upper bounds to accuracy for a given unseen and seen class sets. A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes. A trivial example is if the seen class set is the empty set. A deeper understanding of these requirements may be an avenue to reducing hubness, and therefore improving overall classification performance.
VII Conclusion
A comparison of three generative methods and the baseline method in environmental audio zero-shot learning has been presented. Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly or worse than the baseline, and present higher variance.
The third generative method, ZeroDiffusion, is a novel diffusion-based method for zero-shot learning. It shows a higher accuracy performance than the other three methods, thus producing a new state-of-the-art result in the domain of environmental audio on ESC-50 and FSC22. Diffusion shows potential in zero-shot learning to improve the performance across domains. Future work could include benchmarks for diffusion-based zero-shot learning models on datasets in areas other than environmental audio.
An analysis of confusion matrices and t-SNE graphs of the audio embeddings reveals the impact of audio embedding similarities on the hubness problem and model performance. This analysis is valuable since these similarities may not be intuitive from the raw audio data.
Acknowledgements
This research is supported by an Australian Government Research Training Program Scholarship to the first author.

### v2

@@
-Diffusion in Zero-Shot Learning for Environmental Audio
+Embedding-Space Diffusion for Zero-Shot Environmental Sound Classification
 Abstract
-Zero-shot learning enables models to generalize to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes.
+Zero-shot learning enables models to generalise to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes.
 While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies.
-Generative methods, which have demonstrated success in computer vision, are notably absent from environmental audio zero-shot learning, where classification-based approaches dominate.
+Generative methods, which have demonstrated success in computer vision, are notably absent from zero-shot environmental sound classification studies.
 To address this gap, this work investigates generative methods for zero-shot learning in environmental audio.
 Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN).
-Additionally, a novel diffusion model conditioned on class auxiliary data is introduced.
-The diffusion model generates synthetic data for unseen classes, which is combined with seen-class data to train a classifier.
-Experiments are conducted on two environmental audio datasets, ESC-50 and FSC22.
-Results show that the diffusion model significantly outperforms all baseline methods, achieving more than 25% higher accuracy on the ESC-50 test partition.
-This work establishes the diffusion model as a promising generative approach for zero-shot learning and introduces the first benchmark of generative methods for environmental audio zero-shot learning, providing a foundation for future research in the field.
-Code is provided at https://github.com/ysims/ZeroDiffusion for the novel ZeroDiffusion method.
+Additionally, we introduced a novel diffusion model conditioned on class auxiliary data.
+Synthetic embeddings generated by the diffusion model are combined with seen class embeddings to train a classifier.
+Experiments are conducted on five environmental audio datasets, ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k and TAU Urban Acoustics 2019, and one music classification dataset, GTZAN.
+Results show that the diffusion model outperforms all baseline methods on average across six audio datasets.
+This work establishes the diffusion model as a promising approach for zero-shot learning and introduces the first benchmark of generative methods for zero-shot environmental sound classification, providing a foundation for future research.
 Index Terms:
@@
 Machine learning has revolutionised how society solves complex problems [2, 3, 4].
-Various publications consider the future impact of machine learning, such as AI cities [4], humanoid robotics [5] and medicine [6, 7, 8].
+Various publications consider the future impact of machine learning, such as smart cities [5], humanoid robotics [6] and medicine [7, 8, 9].
 To solve these future problems, machines and AI need to be adaptive [10, 11].
 One way machines can be adaptive is to classify new items without labelled training data [12, 13].
-This problem is called ‘zero-shot learning’.
-Zero-shot learning is a category of supervised machine learning that aims to create a model to classify classes it has not seen during training.
+One method to achieve this is zero-shot learning.
+Zero-shot learning is a subset of supervised machine learning that aims to create a model to classify classes it has not seen during training.
 That is, the set of classes used in training and testing are disjoint.
 Zero-shot learning is valuable in real-world applications, such as remote sensing [14], image classification [15, 16], image segmentation [17, 18], environmental audio [19, 20] and security [21], where there may not exist data for every class.
+Outside of zero-shot learning, recent practical use cases for environmental audio classification include smart cities [22], detection of traffic noise pollution [23], and hearing aid analysis [24].
+These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
 An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem.
@@
 When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings.
-The person performs zero-shot learning using features of the target class and learning what those features mean through knowledge of other animals.
-It is important not to confuse zero-shot learning with similar zero-data problems.
-In zero-shot learning, the method knows about the existence of all classes, i.e.
-the set of classes is closed.
-If it was open, i.e.
-there are unknown classes, the problem is called open-set learning [21, 22, 23].
-Domain adaptation [24, 25] is another problem with a lack of data; however, there is no clear ‘seen’ and ‘unseen’ class distinction.
-An example of domain adaptation is training in simulation and transferring to the real world without fine-tuning [26, 27, 28].
-Finally, zero-shot prompting is a newer area that uses large language models to perform new tasks [29, 30, 31].
-This differs from zero-shot learning as there is no seen and unseen class distinction, and it is hard to quantify how much overlap the task has to scenarios the model was trained on.
-Close to these zero-data problems, particularly zero-shot learning, is few-shot learning.
-Rather than zero data for select classes, few-shot learning has a small number of samples (e.g.
-three to five) for particular target classes.
-A comprehensive review of few-shot learning is provided by Song et al.
-[32].
-In the literature, zero-shot learning methods are commonly tested on computer vision classification tasks [12, 33, 34, 14, 35, 36, 15, 37, 38, 39].
-Traditional methods for zero-shot learning in computer vision relied on classification-based models with engineered attributes [12], but more recent methods use generative models [34, 15, 37] such as generative adversarial networks [40] and variational autoencoders [41].
-These models have been shown to improve performance on unseen classes and can help mitigate problems such as the hubness problem and the domain shift problem [42].
-The hubness problem occurs when data points are concentrated around a minority of classes, leading to poor performance when tested on unseen classes [43].
-The domain shift problem arises when the model performs poorly because the training and test sets have different distributions [44].
-This work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision.
+The person performs zero-shot learning using attributes of the target class and learning what those attributes mean through knowledge of other animals.
+The presented work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision.
 It takes inspiration from the successful usage of generative methods in computer vision [15, 25, 26, 27], and from a newer generative method, diffusion, which has seen success in image generation tasks [28, 29].
 There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
-Recent use cases for environmental audio classification include smart cities [48], detection of traffic noise pollution [49], and hearing aid analysis [50].
-These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
-There is a lack of datasets that can be used to benchmark environmental audio zero-shot learning, as datasets are often not released in their raw audio form or lack classes.
-AudioSet [51] is a valuable large dataset, but instead of audio clips, it contains embeddings.
-Embeddings are high dimensional vectors representing the data.
-These embeddings come from a network pre-trained on AudioSet.
-This pre-training makes the dataset unsuitable for zero-shot learning, as the unseen classes may be present.
-Other datasets, such as UrbanSound8k [52], have too few classes for rigorous zero-shot learning.
-Partitioning seen and unseen classes will result in a seen set with little diversity and may not cover a large enough semantic space for the model to generalise to unseen classes, risking overfitting.
-Additionally, the unseen classes will be very few, and so the statistical significance of the results is poor.
-A reasonable minimum may be five unseen classes, and double that amount of seen classes, for a total of 15 classes in the dataset.
-ESC-50 [53] is an open dataset with fifty classes, often used in environmental audio zero-shot learning literature.
-FSC22 [54] is another open dataset and contains twenty-seven classes.
-While this is almost half the size of ESC-50, it is still a substantial size to perform zero-shot learning on.
-We are not aware of the existence of any zero-shot learning literature using this dataset.
-The present study uses ESC-50 and FSC22 to compare three generative methods.
-These methods use the following architectures: generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion networks.
-A bilinear compatibility network [55] is also compared as a baseline.
-The method ZeroDiffusion is introduced as a new method for zero-shot learning using diffusion.
-The GAN method used is LisGAN [15]; however, the component where high-confidence unseen samples are injected back into the training is not included, to prevent violating the zero-shot learning problem formation where unseen class data should not be present during training.
-The method uses a conditional Wasserstein GAN [56].
-Zero-shot learning methods often aim to learn the relationship between feature and class embeddings.
-To achieve this, the GAN is conditioned on class embeddings.
-It is then used to generate synthetic samples for unseen classes.
-Vanilla conditional GAN methods were experimented with in the presented study, but they suffered extreme instability and low performance.
-The adjusted LisGAN method performs worse than the bilinear compatibility method but is better than a random model.
-The cross-aligned and distribution-aligned VAE (CADA-VAE) [14] has been previously used in computer vision with success.
-The method trains two variational autoencoders, one for auxiliary data and one for feature data, to learn a shared latent space.
-The results were comparable to the bilinear compatibility method.
-Finally, the most successful model is our new diffusion model.
-The experiments with diffusion show a clear improvement upon any previous method in environmental audio zero-shot learning.
-While there is existing zero-shot learning literature outside of environmental audio on diffusion models [57, 58], no method trains a network from scratch.
-The methods use the pre-trained Stable Diffusion model [47].
-While this can provide practical usage, it is hard to draw theoretical conclusions on the zero-shot performance of this method.
-The overlap of the unseen classes and training classes is unknown.
-This work shows that diffusion, with a transparent training process and a valid seen and unseen class distinction, outperforms other techniques in zero-shot learning.
+Six datasets - ESC-50, FSC22, ARCA23K-FSD, TAU Urban Acoustic Scenes 2019, UrbanSound8k and GTZAN - are used to compare four methods - the baseline non-linear compatibility network [30] based on Attribute Label Embedding (ALE) [31], LisGAN [16], CADA-VAE [15] and our new novel method, ZeroDiffusion.
+This represents a broad study of generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion for zero-shot learning, with the baseline as a comparison.
+ZeroDiffusion achieves significant accuracy improvement in most cases.
 The contributions of the presented paper are:
-  • The first diffusion method for zero-shot learning that does not use a pre-trained model.
-  • A new state-of-the-art for zero-shot learning with the ESC-50 dataset using the diffusion method.
-  • The first benchmark on the FSC22 dataset, with the diffusion method outperforming other methods.
-  • The first comparison of GANs, VAEs and diffusion-based methods for environmental audio zero-shot learning with ESC-50 and FSC22.
+  • The first diffusion method for zero-shot learning that does not use a pre-trained model, but rather works in the embedding space.
+  • A new state-of-the-art method for zero-shot environmental sound classification that performs equally to or outperforms a range of existing methods in zero-shot learning across six audio datasets.
+  • The first zero-shot learning benchmark on the FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019 datasets.
+  • The first comparison of GANs, VAEs and diffusion-based methods for zero-shot environmental sound classification with six audio datasets: ESC-50, FSC22, ARCA23K-FSD, UrbanSound8k, TAU Urban Acoustic Scenes 2019 and GTZAN.
 II Related Work
@@
 Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision.
-Finally, existing works in environmental audio zero-shot learning are considered.
+Next, existing works in zero-shot environmental audio classification are considered and audio datasets are covered.
+Finally, with this context, the novelty and contributions of the presented work are explained.
 II-A Terminology
-In this section, zero-shot learning terminology is presented in an effort to contextualise the literature.
-First, embeddings for both class and feature data are considered, then the concept of auxiliary data and finally, a high-level definition of zero-shot learning.
+To contextualise the literature, this subsection introduces foundational terminology in zero-shot learning.
+It first describes class and feature embeddings, followed by the role of auxiliary data.
+A high-level definition of zero-shot learning is then provided to frame the broader research landscape.
 II-A1 Embeddings
-Embeddings are a method of converting raw data, such as images or audio, into a high-level form that machine learning algorithms can use.
-The form embeddings take is a high-dimensional vector of real numbers.
-The methods in this work use deep learning to obtain these vector representations, using both supervised and unsupervised methods.
+Embeddings convert raw data, such as images or audio, into a high-level form that machine learning algorithms can use.
+The form embeddings take is a high-dimensional vector of real numbers, formally $x\in\mathbb{R}^{\alpha}$ is an example of an embedding with dimension $\alpha$.
+The methods in this work use deep neural networks to obtain these vector representations, using both supervised and unsupervised methods.
 An example of what an embedding space might look like is shown in Figure 1.
-Lions and cats are both felids and therefore semantically similar and closer to each other in the vector space.
-‘Dog’ is semantically closer to the felids than ‘car’ as they are types of animals.
+Lions and cats are semantically similar and, therefore, are closer to each other in the vector space than a dog or car.
 In this way, the structure of the space reveals semantic relationships.
 Figure 1: An embedding space. Semantically similar data is closer, such as ‘lion’ and ‘cat’.
-Supervised methods to create embeddings typically involve classical supervised learning on a dataset and the output of the second last layer of the artificial neural network becomes the embedded data vector.
+Supervised methods to create embeddings typically involve classical supervised learning on a dataset, and an internal layer of the artificial neural network becomes the embedded data vector.
+In many cases, the second last layer is used because it is expected to contain high-level meaningful information compared to previous layers.
 Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set.
-Typically the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset could be used as long as there is no overlap with the unseen class set.
-Unsupervised methods involve learning a mapping to a space where the semantic distance of data is represented in the space.
-The most common example of this embedding method is word embeddings, such as GloVe [59] and Word2Vec [60].
+Typically, the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset is suitable as long as there is no overlap with the unseen class set.
+Unsupervised methods involve learning a mapping to a space that represents the semantic distance of data.
+A common example of this type of embedding method is word embeddings, such as GloVe [32] and Word2Vec [33].
 These methods learn the semantic similarity of words using their proximity in texts.
@@
 For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [33].
+GloVe uses matrix factorisation and builds a large co-occurrence matrix of word pairs to globally understand the relationship between words, while Word2Vec uses context-based predictive neural networks.
+While their learning processes are different, both methods have the same goal and provide semantically rich embeddings.
 II-A2 Auxiliary Data
 Zero-shot learning involves learning to classify known classes without samples for those classes.
-If the classes are unknown, the problem is called open-set learning.
+If the classes are unknown, the problem is open-set learning.
 These known classes contain information even without samples.
-This information is called ‘auxiliary data’.
+This information is ‘auxiliary data’.
 The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe.
-Another popular form is human-engineered attributes, where a vector is created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’.
-The Animals with Attributes [12] and Animals with Attributes 2 [61] datasets are examples of datasets that contain human-engineered attribute vectors, and were created for zero-shot learning research.
+Another popular form is human-engineered attributes, where a vector is manually created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’.
+The Animals with Attributes [13] and Animals with Attributes 2 [34] datasets are examples of datasets that contain human-engineered attribute vectors and were created for zero-shot learning research.
 II-A3 Zero-Shot Learning
-Zero-shot learning involves training on seen classes and testing on unseen classes.
-Generalised zero-shot learning extends this problem to include both seen and unseen classes in testing.
-Zero-shot learning is one of many zero-data problems, as described in the introduction.
-Feature data is defined as the target data the user wants to classify.
-In computer vision, the feature data is images.
-In environmental audio, it is audio clips.
-Feature data for unseen classes should not be present anywhere before the testing phase, including during pre-training.
-The embeddings for the feature data and classes represent semantic information.
-Zero-shot learning aims to train a model that generalises its understanding of semantics to go beyond class-specific knowledge.
+Zero-shot learning involves training a model on a set of seen classes $X$ and evaluating its performance on unseen classes $Y$, where $X\cap Y=\varnothing$.
+Feature data refers to the input the model is classifying, such as images in computer vision or audio clips in environmental sound classification.
+Feature data from unseen classes must not be used before testing, including pre-training.
+Both feature data and class labels are represented as embeddings that capture semantic information.
+The goal of zero-shot learning is to train a model that can leverage these semantic representations to generalise beyond the classes it has been explicitly trained on.
 II-B Generative Methods in Zero-Shot Learning
@@
 These largely involve networks combining GANs and autoencoders.
+II-B1 Autoencoders
 In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al.
@@
 [36].
-These used two traditional autoencoders for each modality and learnt a shared latent space between these two autoencoders.
+They used two traditional autoencoders for each modality (class textual descriptions and images) and trained a shared latent space between these two modalities.
 Conditional variational autoencoders were introduced in zero-shot learning the following year [26].
-A new benchmark was found when CADA-VAE [14] was proposed which took advantage of variational autoencoders and created a shared latent space through cross-alignment and distribution-alignment.
+A new benchmark was found when CADA-VAE [15] was proposed, which used variational autoencoders and strengthened the shared latent space through cross-alignment and distribution-alignment of the two modalities.
+II-B2 Generative Adversarial Networks
 One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [37].
-In the same year, another GAN method was created that used a visual pivot regulariser to guide training [63].
-It also used Wikipedia text descriptions converted to word embeddings using term frequency-inverse document frequency (TF-IDF).
-The method was compared with other zero-shot learning methods including a basic GAN model.
-The basic GAN model performs much worse than any established zero-shot learning method, while Zhu et al.’s method [63] shows an improvement over the state-of-the-art.
-LisGAN [15] improved on previous GAN methods by using a Wasserstein GAN conditioned with auxiliary data.
-They use ‘soul samples’, samples representing the ‘ideal’ sample from a class.
-Soul samples for seen classes are the average sample over all the seen samples for that class, while soul samples for unseen classes are constructed using the auxiliary data for the class.
-The GAN is trained to generate data close to a soul sample.
-The authors perform a step where unseen data is tested and any samples with high confidence are injected into training.
-This violates the requirement in zero-shot learning where unseen data should not be seen during training, even if it is unlabelled.
+A Wasserstein GAN is a vanilla GAN with the loss function changed to measure the Wasserstein distance between the real and generated data distributions [38].
+It results in more stable training and mitigates mode collapse.
+In the same year, Zhu et al.
+introduced another GAN method using a visual pivot regulariser to guide training [39].
+The regulariser encourages the generation of features for a class to be close to a fixed pivot point.
+It used Wikipedia text descriptions converted to word embeddings using the term frequency-inverse document frequency (TF-IDF) as class data.
+The authors compare their method with other zero-shot learning methods, including a basic GAN model.
+The basic GAN model has much worse accuracy than any established zero-shot learning method on the vision classification task across benchmark datasets.
+LisGAN [16] improved on previous GAN methods by using a Wasserstein GAN conditioned on auxiliary data.
+They use ‘soul samples’, samples representing the ‘ideal’ sample from a class, to guide generation.
+The method includes a self-training step where high-confidence samples from the unseen classes are injected into the training process.
+This makes the approach transductive, since it assumes access to unlabelled unseen data, in contrast to the inductive setting commonly assumed in zero-shot learning where no unseen data is available.
 GDAN [27] proposes a GAN with a regressor and cyclic-consistent loss.
 The results are marginally better than the conditional VAE method from Mishra et al.
-[34], indicating that GAN methods still fall behind VAE methods for zero-shot learning, given that at this point better VAE methods had been introduced.
+[26].
+Given the existence of superior VAE methods, this indicates that GAN methods still fall behind VAE methods for zero-shot learning.
 Task Aligned Generative Meta-learning for Zero-shot Learning [40] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes.
-The backbone of the technique is a meta conditional GAN [64].
+The backbone of the technique is a meta conditional GAN [41], which combines a meta-learning network and a conditional GAN.
 SAGAN [42] involves training a GAN, followed by training a classifier to align the visual and semantic space.
-This method considers what to do with the trained GAN rather than improving the GAN training method.
-In the same year, ZeroNAS [65] is proposed which performs neural architecture search on GAN zero-shot learning methods to find the optimal architecture for a zero-shot learning problem.
-The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of the expert curation of these networks.
-While there is much literature on how GANs may be used for zero-shot learning and how general improvements to GAN training can be used, they are still notoriously hard to train [63, 65].
-This does pose the question of if GANs are too unstable and not reliable enough for zero-shot learning problems.
+This method focuses on improving the classifier rather than the GAN.
+In the same year, ZeroNAS [43] is proposed which performs neural architecture search on GAN-based zero-shot learning methods to find the optimal architecture for a zero-shot learning problem.
+The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of expert curation of these networks.
+While there is much literature on using GANs in zero-shot learning, they are still notoriously hard to train [39, 43].
+GANs may be too unstable and unreliable for zero-shot learning problems.
 In practical situations, it is hard to estimate model performance on unseen classes.
 Therefore, the method must be reliable.
-Chou et al.
-[35] propose a generative method without VAEs or GANs.
-It uses a beta distribution to generate virtual classes for more robust zero-shot learning.
-Despite their efforts, GANs and VAEs still dominate the landscape of zero-shot learning.
+II-B3 Diffusion
 In 2023, two publications emerged using diffusion in zero-shot learning [44, 45].
@@
 These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning.
+Additionally, since Stable Diffusion is foundational to the method, usage is limited to problems in computer vision and the use of textual class auxiliary data.
 II-C Environmental Audio Zero-Shot Learning
 Limited research exists for environmental audio zero-shot learning.
-One of the earliest works proposes a Twin neural network [66], tested on a small dataset containing ten classes.
-Xie and Viranen [18] used a more substantial dataset of fifty classes.
-In addition to a bilinear compatibility model [67] that created a baseline for the ESC-50 dataset, they proposed two partition strategies for five-fold cross-validation.
+An early study proposes a Twin neural network [46], tested on a small dataset containing ten classes.
+A Twin neural network is an architecture with two identical networks that process two inputs at the same time to learn a similarity function.
+Xie and Virtanen [19] used a more substantial dataset of fifty classes.
+They introduce a bilinear compatibility network, which builds on from Attribute Label Embedding (ALE) [31].
+A compatibility function is learnt between the audio data and class embeddings by computing:
+$f(x,y)=x^{T}Wy$ (1)
+Where $x$ is the audio embedding, $y$ is the class embedding, and $W$ is a trainable weight matrix.
+This maps the inputs to a shared latent space for comparison directly, without the use of generative methods.
+This method sets an initial baseline for the ESC-50 dataset, and the authors introduce two partition strategies for five-fold cross-validation.
 These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
 Further work in the field [47, 48] brought new state-of-the-art results and critical considerations when training zero-shot networks.
-Xie and Viranen [68] highlight the importance of the unseen and seen class distinction, where they identify that VGGish pre-trained on AudioSet is unsuitable for zero-shot learning on ESC-50 due to the presence of unseen classes in AudioSet.
-AudioSet contains embeddings rather than audio clips.
-Embeddings are a critical aspect of zero-shot learning methods, and some research works explore improving these embeddings [19].
-Furthermore, it is vital to maintain the distinction between ‘seen’ and ‘unseen’ classes.
-Therefore, training from raw data is essential.
-AudioSet provides the location of the YouTube videos the clips come from, but it is against the Terms of Service to download the raw data.
-AudioSet is, therefore, unsuitable for zero-shot learning research.
+Xie et al.
+[47] highlight the importance of the unseen and seen class distinction when using pre-trained models.
+A VGGish model trained on AudioSet (a large environmental audio dataset) may produce rich embeddings, but a majority of classes in ESC-50 occur in AudioSet.
+The classes in ESC-50 which are meant to be unseen would have instead been seen by the VGGish network during pretraining on AudioSet.
+This breaks the zero-shot learning formulation.
+It is important to understand that it is the feature data that should be unseen - the classes are known, and therefore a pre-trained word embedding model does not invalidate the method as zero-shot learning.
+Wu et al.
+use contrastive audio-text representation learning with CLAP to perform zero-shot learning on multiple classes[49].
+However, like other CLAP and CLIP-based methods, these models require training data from large datasets in the target modality and this is not always feasible in a given domain.
+Additionally, even though repeated samples across datasets are removed, whole classes aren’t removed.
+This is evident in the ESC-50 and FSD50k case, where the text reports 630 overlapping files, even though within one overlapping class, ‘dog’, there are over 900 samples in FSD50K that would need to be removed.
+Indeed, this would result in extremely few classes remaining since audio sources for datasets are very limited and many datasets are from YouTube or Freesound, with many of the same classes appearing across audio datasets.
+However, this inflates the reported zero-shot classification performance.
 Some research focuses on improvements through embeddings.
-While not often the main focus of methods, they are the foundations of the zero-shot method.
-The remaining pipeline is limited to the information that the embeddings can capture.
+While not often the main focus of methods, they provide the foundations for unseen classification.
+Any embedding-based method is limited to the information that the embeddings can capture.
 The methods in this paper use work by Sims et al.
-[19] to improve the performance through embeddings by adding synonyms.
-III Method
+[20] to improve the performance through embeddings by adding synonyms, except ARCA23K-FSD (due to size) and GTZAN (lack of synonyms for genres).
+II-D Contributions and Novelty
+This section will detail the contributions of the presented paper, in the context of the literature.
+Literature for zero-shot environmental sound classification is lacking, and no existing literature in the domain explores generative methods.
+The experiments in this work not only benchmark environmental audio datasets on the existing generative methods CADA-VAE and LisGAN but also introduce a novel diffusion-based method that is not tied to a particular modality.
+Code for the novel ZeroDiffusion method is provided at https://github.com/ysims/ZeroDiffusion.
+Two publications exist in zero-shot learning using diffusion ([44, 45]); however, they both use Stable Diffusion directly to classify unseen data, with no further training of any model.
+Their methods are specific to computer vision and textual class labels through their use of Stable Diffusion.
+Meanwhile, the novel diffusion method presented in this work is modality-agnostic, because it uses embeddings as input to the diffusion model.
+Any class embedding can be used to condition the diffusion model, and any feature embedding can be used for training, regardless of what modality it comes from.
+The method involves training a diffusion model, which is used to generate unseen embeddings to train a compatibility-based classifier.
+This generative-style zero-shot learning method is significantly different to the previous work on Stable Diffusion and provides greater flexibility.
+In addition to designing the first modality-agnostic diffusion method for zero-shot learning, this work additionally performs the first zero-shot benchmarks of FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019.
+III Methods
 This section will begin with a technical overview of zero-shot learning to complement the high-level background given in the previous section.
@@
 III-A Zero-Shot Learning
-The zero-shot learning problem is defined as the following.
-There exists a set $X$ of seen data and a set $Y$ of unseen data such that $X\cap Y=\varnothing$.
-The zero-shot learning method aims to learn a function
-$f(z,C)=c$ (1)
-where $z\in X\cup Y$ is a data sample; $C$ is the set of all classes; and $c\in C$ is the class that the sample $z$ comes from.
-After training a model to learn $f$ using all samples in $X$, the model is tested on all samples in $Y$.
-The literature frequently divides zero-shot learning methods into three components.
-One component is the auxiliary data function $\sigma(c)$ that takes a class $c\in C$.
+The zero-shot learning method aims to learn a similarity function $f(w,z)$, where $w$ is a feature data sample and $z\in X\cup Y$ is a class.
+$X$ is the set of unseen classes and $Y$ is the set of seen classes.
+$X\cup Y$ is the set of all classes and $X\cap Y=\varnothing$.
+After training a model to learn $f(w,z)$ using all samples belonging to classes in $X$, the model is tested on all samples belonging to classes in $Y$.
+Many zero-shot learning methods can be considered to contain three distinct components.
+One component is the auxiliary data function $\sigma(z)$ that takes a class $z\in X\cup Y$.
 The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$.
 $\sigma$ changes a class label into high-level information.
-A second component is the feature embedding function $\tau(z)$, where $z\in X\cup Y$ is any feature data in the dataset.
+A second component is the feature embedding function $\tau(w)$, where $w$ is any feature data in the dataset.
 The feature data in the case of environment audio is audio clips.
@@
 The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer.
-In computer vision, a popular method is to use ResNet, pre-trained on ImageNet.
-The third component ties the previous two together to retrieve a class prediction. This component learns a function
-$\xi(\tau(z),\sigma(C))=c$ (2)
-that takes the high-level information output of the other two components and outputs the corresponding class $c$ for a sample $z$.
-This component of zero-shot learning is explored heavily across the literature.
+The third component ties the previous two together to retrieve a class prediction.
+This component learns a similarity function $\xi(\tau(w),\sigma(z))$ that takes the high-level output of the two embedding components and outputs the similarity of $w$ and $z$.
+By comparing $w$ against all classes ($X$ during training and $Y$ during testing), either top-1 or thresholding can be used to retrieve the prediction.
 III-B Embeddings
-An embedding is an encoded form of data that is easy for a machine learning model to process.
+An embedding is an encoded form of data that is simple for a machine learning model to process.
 They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
@@
 Learning the relationship between the feature and class semantic embeddings is crucial to performing zero-shot learning.
-The $\sigma$ function is the Word2Vec [60] pre-trained model.
+The $\sigma(z)$ function in this work is the Word2Vec [33] pre-trained model.
 This model takes class labels and outputs an embedding vector for that word.
@@
 For ESC-50, the synonyms from the original paper were used.
-FSC22 has not been used in previous literature on synonyms, and so we human-engineer the synonyms, where overlapping classes use the same synonym list from ESC-50.
-To learn $\tau$ and obtain audio embeddings, we train a modified YAMNet model, shown to perform better than other methods [19], on the seen class data.
+Synonyms are human-engineered for the other datasets, where overlapping classes across datasets use the same synonym lists.
+ACRA23K-FSD and GTZAN are used without added synonyms, due to size and difficulty in getting synonyms for music genres.
+Synonyms used are detailed in the supplementary materials.
+To learn $\tau(w)$ and obtain audio embeddings, we train a modified YAMNet model [20] on the seen class data, due to the higher accuracy compared to VGGish when used for zero-shot learning.
+The input to the modified YAMNet model is a log Mel spectrogram with 64 bins.
 As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
-Thus both $\sigma$ and $\tau$ have been defined, where the output of $\sigma(c)$ is a class embedding of dimension $300$ for the class $c$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w\in X\cup Y$.
+Both $\sigma(z)$ and $\tau(w)$ have been defined, where the output of $\sigma(z)$ is a class embedding of dimension $300$ for the class $z\in X\cup Y$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w$.
 III-C ZeroDiffusion
 Figure 2: ZeroDiffusion involves training three models.
-The first is the audio embedding model, used to retrieve audio embeddings.
-Seen classes are used to train this model.
-Next, the diffusion model is trained using audio embeddings from the first model and word embeddings.
-This process is done using seen classes.
-Finally, a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
+The first (1) is the audio embedding model, used to retrieve audio embeddings.
+The model is trained with seen classes.
+Next, (2) the diffusion model is trained using audio embeddings and word embeddings from seen classes.
+Finally, (3) a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
+We introduce our new method ZeroDiffusion, which applies a novel embedding-space diffusion model to generate synthetic unseen data to train a compatibility classifier.
+While other diffusion methods use large diffusion models to generate feature embeddings, ZeroDiffusion uses feature embeddings as input to learn to generate unseen data.
+The diffusion model itself is very lightweight because it works within the embedding space.
 Diffusion involves training a model to reconstruct data from noise.
-During training, the input to the model is a data sample with noise added, and the expected output is the original data sample.
+During training, the model input is a data sample with added noise, and the expected output is the original data sample.
 The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise.
@@
 If a user requests output of ‘a cat holding a basketball’, the model should be able to recreate a picture of a cat holding a basketball from noise.
-A diffusion model conditioned on auxiliary data is trained with the input to the model being paired embeddings $\sigma(c)$ and $\tau(w)$ where $c$ is the corresponding class for the audio sample $w\in X$.
-The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(c)$.
-The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with mean 0 and variance 0.1 to create a gradually noisy embedding vector
-$\tilde{\tau}(w)(p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
-where $p$ is the epoch percentage.
-The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(c)$.
+A diffusion model conditioned on auxiliary data is trained where the input to the model is the paired embeddings $\sigma(z)$ and $\tau(w)$ where $z$ is the corresponding class for the audio sample $w$.
+The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(z)$.
+The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with a mean of 0 and a variance of 0.1 to create a gradually noisy embedding vector:
+$\tilde{\tau}(w,p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (2)
+where $p$ is the epoch percentage. Similarly the class embedding $\sigma(z)$ is jittered for diversity using
+$\tilde{\sigma}(z)=\sigma(z)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
+The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(z)$.
 The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428.
@@
 The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size.
-This is followed by batch normalisation for training stabilisation and then a Leaky ReLU activation to add a non-linearity component.
-Following Leaky ReLU, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space.
+This is followed by a Leaky ReLU activation for a non-linearity component and then a dropout of $0.3$ to encourage diversity.
+Following the dropout layer, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space.
 The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
-The loss is the mean squared error between the embedding the model is trying to recreate and the output of the model.
+The loss term involves five components: a mean squared error for the reconstruction loss, a maximum mean discrepancy for distribution loss, variance loss to encourage diversity, an average mean squared error of generated and real features in a batch for global centroid alignment, and a cosine similarity loss to encourage directional alignment of embeddings.
 During training, gradients are clipped to a maximum norm of 1 to stabilise training.
@@
 Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise.
-After creating an unseen dataset in this manner, a bilinear compatibility network [67, 18] is trained with both the true seen data and synthetic unseen data.
-The bilinear network uses a cross-entropy loss function and an Adam optimiser.
+After creating an unseen dataset in this manner, a non-linear compatibility network [31, 48] is trained with both the true seen data and synthetic unseen data.
+The compatibility network uses a cross-entropy loss function and an Adam optimiser.
+III-D Non-linear ALE
+Xie et al.
+introduce the non-linear version [48] of the bilinear compatibility network [19], which is based on ALE [31].
+This method is used as a baseline due to its good performance amongst the zero-shot environmental sound classification literature and its simplicity.
+The lack of generators (as opposed to GAN and VAE methods) and its single classification nature means that it has few hyperparameters to tune and training across runs is stable, as shown in the experiments presented in section IV.
+The method learns a function:
+$f(w,z)=(B\cdot(\tanh(A\cdot w)))^{T}\cdot z$ (4)
+where $w$ is the audio embedding, $z$ is the class embedding, $A$ and $B$ are both trainable weight matrices, and $\tanh$ is the hyperbolic tangent activation.
+$f(w,z)$ represents the similarity between $w$ and $z$, and is calculated for every $z$.
+This creates an encoded vector where the $i$th position in the vector represents the confidence or similarity for the $i$th class.
+The model is independent of the number of classes, unlike typical classification models where the last layer is equal to the number of classes.
+III-E LisGAN
+LisGAN [16] is a GAN method that uses ‘soul samples’ to regulate data generation.
+It trains a conditional Wasserstein GAN [38] where the generator uses class embeddings and noise to generate feature data and the discriminator not only determines whether a sample is generated or real but also has a supervised classification component to guide the conditional training.
+Each class has a soul sample - a representative sample for the class - and the generator is trained to generate samples for a given class close to that class’s soul sample.
+For seen classes, the soul sample is the average of that class’s data.
+For unseen classes, the class embedding defines the soul sample.
+The discriminator’s classification output is used to screen synthetic samples.
+Synthetic samples with high confidence are used to train a final classifier.
+Additionally, unseen and unlabelled data that the discriminator’s classifier predicts with high confidence are used in training the final classifier.
+This use of unlabelled unseen data is transductive zero-shot learning; however, in this work, only the inductive case is considered.
+Therefore, we remove the training on unlabelled seen data.
+The core idea of the LisGAN method remains, where soul samples are used to guide GAN training, and the discriminator’s classifier is used to filter data to train a final classifier.
+Despite removing a component of LisGAN, it outperformed other GAN methods (vanilla conditional Wasserstein GAN, f-CLSWGAN) in pilot studies.
+III-F CADA-VAE
+CADA-VAE [15] is a popular method in computer vision-based zero-shot learning that competes with new methods.
+In this work, CADA-VAE is applied to the domain of environmental audio to evaluate its performance as a variational autoencoder-based method for zero-shot environmental sound classification.
+The core of the method is the two variational autoencoders (VAEs) and the loss function.
+One VAE is designed to encode and decode class embeddings, while the other is designed to encode and decode feature embeddings (typically image embeddings, but in this work, audio embeddings are used).
+Not only are the VAEs trained in a typical manner with reconstruction loss and Kullback–Leibler divergence, but they are also trained with cross-alignment and distribution alignment.
+Cross-alignment measures the ability of the decoders to decode data from the other modality into their modality - that is, take an encoded class embedding and decode it into a feature embedding, and vice versa.
+Distribution alignment measures the difference between the distribution of the class embedding autoencoder and the feature embedding autoencoder.
+Both of these loss terms aim to create a shared latent space between both modalities.
+The latent space is then used to train a classifier, which performs zero-shot learning at test time.
 IV Experiments
-Experiments are performed on the ESC-50 and FSC22 datasets.
-The state-of-the-art in zero-shot learning environmental audio, the compatibility bilinear method [67, 18] is compared with existing methods from computer vision [14, 15], and our new method ZeroDiffusion.
-ESC-50 is split into five partitions of ten classes.
-One partition is held as the test set.
-The remaining four partitions are used for four-fold cross-validation, where three of the partitions are used in training and the remaining one is used for validation.
-The validation set is used to tune the model and its hyperparameters before using the test set.
-This design is modified from the five-fold cross validation used in previous work [68, 19] to include a held-out test set.
-Using a four-fold strategy for hyperparameter tuning allows for more generalised hyperparameter tuning and a better understanding of the performance of the model.
-The final experiments on the held-out test set give a good indication of how the model performs on a completely unseen set of classes and data.
-ESC-50 is a suitable dataset for this strategy as it is composed of many classes, allowing for each partition to have ten classes.
-The dataset is correctly labelled and clean, making it ideal for testing.
-FSC22 is split into three sets, a test set with 7 classes, a validation set of 7 classes, and a training set with the remaining 13 classes.
-The validation set is used for hyperparameter tuning, and then the method is evaluated using the test set.
-This dataset has almost half the number of classes compared to ESC-50, reducing the amount of flexibility in forming a strategy.
-Additionally, the number of classes is uneven and has few divisors.
-If it were partitioned into nine lots of three, the number of classes in each set would be extremely small for zero-shot learning while also creating a lot of folds to manage.
-If three lots of nine were used, the training set would be very small compared to the usual split in zero-shot learning, where there are more training classes than test classes.
-Therefore, the validation and test sets are chosen to be the same size for consistency, and train with a larger number, being almost the size of the test and validation sets combined.
-Therefore, $k$-fold validation is not performed for this dataset.
-The dataset does overlap with the classes in ESC-50, but there are new classes not present in ESC-50.
-Some classes have questionable labels, including the class ‘silence’, which appears to contain a variety of soundscapes including birds chirping and rain falling, which are classes elsewhere in the dataset.
-Because the dataset is not entirely well-labelled, this is considered a harder dataset compared to ESC-50.
-All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were run on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
-Index Class Index Class Index Class
-0 Dog 17 Pouring water 34 Can opening
-1 Rooster 18 Toilet flush 35 Washing machine
-2 Pig 19 Thunderstorm 36 Vacuum cleaner
-3 Cow 20 Crying baby 37 Clock alarm
-4 Frog 21 Sneezing 38 Clock tick
-5 Cat 22 Clapping 39 Glass breaking
-6 Hen 23 Breathing 40 Helicopter
-7 Insects (flying) 24 Coughing 41 Chainsaw
-8 Sheep 25 Footsteps 42 Siren
-9 Crow 26 Laughing 43 Car horn
-10 Rain 27 Brushing teeth 44 Engine
-11 Sea waves 28 Snoring 45 Train
-12 Crackling fire 29 Drinking, sipping 46 Church bells
-13 Crickets 30 Door knock 47 Airplane
-14 Chirping birds 31 Mouse click 48 Fireworks
-15 Water drops 32 Keyboard typing 49 Hand saw
-16 Wind 33 Door, wood creaks
-TABLE I: Class indices and names in ESC-50. This follows the convention from the dataset.
-Index Class Index Class Index Class
-0 Fire 9 Axe 18 Footsteps
-1 Rain 10 Chainsaw 19 Clapping
-2 Thunderstorm 11 Generator 20 Insect
-3 Water Drops 12 Handsaw 21 Frog
-4 Wind 13 Firework 22 Bird Chirping
-5 Silence 14 Gunshot 23 Wing Flapping
-6 Tree Falling 15 Wood Chop 24 Lion
-7 Helicopter 16 Whistling 25 Wolf Howl
-8 Vehicle Engine 17 Speaking 26 Squirrel
-TABLE II: Class indices and names in FSC22. While the same order is followed from the dataset, here the classes are zero-indexed rather than one-indexed.
+Experiments are performed on the ESC-50, ARCA23K-FSD, FSC22, UrbanSound8K, TAU Urban Acoustic Scenes 2019 and GTZAN datasets.
+The state-of-the-art in zero-shot learning environmental audio, the non-linear ALE method [31, 48] is compared with existing methods from computer vision [15, 16], and our new method ZeroDiffusion.
+All experiments are repeated ten times with the average and standard deviation of the accuracy recorded.
+The experiments were conducted on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
+IV-A Datasets
+Environmental sound classification severely lacks large datasets and diversity.
+ESC-50, FSC22, ARCA23K-FSD and UrbanSound8k are sourced from Freesound, a repository of user-contributed tagged audio clips released under the Creative Commons license.
+Some clips in these datasets may overlap, and classes are typically similar.
+TAU 2019 and GTZAN are collected from other sources, and investigate the problems of acoustic scene recognition and music genre classification, respectively.
+These datasets add diversity to the experiments both in audio clip source and task.
+IV-A1 ESC-50
+ESC-50 [51] is a clean environmental sound classification dataset of 50 classes which can be arranged into five balanced categories with ten classes each.
+There are 40 audio clips per class, each five seconds long.
+The dataset is from Freesound and has been carefully curated to ensure a clean dataset.
+It is a popular dataset due to its reasonable size and correct labels.
+The large number of classes in the dataset motivates using k-fold cross-validation, as done in previous work [19, 20].
+IV-A2 FSC22
+FSC22 [52] is a sound classification dataset focusing on forest sounds.
+It has 27 classes, each with 75 5-second-long audio clips.
+The clips are sampled from Freesound.
+The dataset is not as clean as ESC-50, particularly the ‘silence’ class where some samples contain the sound of bird calls, despite the existence of a bird class.
+Because of this, the dataset is considered harder to learn.
+IV-A3 ARCA23K-FSD
+ARCA23K-FSD [53] is a single-label dataset for environmental audio classification derived from the multi-label FSD50K environmental audio dataset, which in turn is collected from Freesound.
+The dataset contains 70 unbalanced classes and over 23,000 clips.
+FSD50K is manually human-labelled and contains clips varying in length from 0.3 seconds to 30 seconds.
+Due to the large number of classes, a similar approach to ESC-50 can be undertaken where k-fold cross-validation can be used for evaluation.
+IV-A4 UrbanSound8k
+UrbanSound8k [54] is an environmental sound classification dataset collected from Freesound.
+Each clip is no more than 4 seconds long, and a total of 8,732 clips are in the dataset.
+There are 10 unbalanced classes.
+The small amount of classes is problematic for zero-shot learning, as even splitting into three sets (train, validation and test) results in very little in each partition.
+An alternative is to use only train and test partitions.
+IV-A5 TAU Urban Acoustic Scenes 2019
+TAU 2019 [55] is an acoustic scene dataset where the aim isn’t to classify specific sounds but rather the type of location the sound is from, for example, a park or an airport.
+Each audio segment is 10 seconds long and belongs to one of 10 acoustic scene classes.
+Each scene has 1,440 audio clips.
+The audio clips in the dataset were collected by the authors rather than scraped or curated from existing sources online.
+This positions it as a useful dataset due to its unique audio clips, whereas the Freesound datasets have overlap due to their shared source.
+Like UrbanSound8k, there are only 10 classes so splitting the data into two sets (train and test) is more reasonable than adding a validation partition.
+IV-A6 GTZAN
+GTZAN [56] is a music genre dataset of 10 genres with 100 audio files for each class.
+Each clip is 30 seconds long.
+While this is an adjacent domain to environmental sound classification, it provides diversity to the experiments when many of the other datasets come from the same source (Freesound).
+The number of classes is equal to TAU 2019 and UrbanSound8k, motivating the use of the same partitioning strategy.
 Partition Classes
@@
 fold 4 0, 5, 8, 9, 11, 12, 15, 16, 37, 43
+ARCA23K-FSD fold 0 crash cymbal, run, zipper (clothing), acoustic guitar, gong, knock, train, crack, cough, cricket
+fold 1 electric guitar, chewing mastication, keys jangling, female speech woman speaking, crumpling crinkling, skateboard, computer keyboard, bass guitar, stream, toilet flush
+fold 2 tap, water tap faucet, squeak, snare drum, finger snapping, walk footsteps, meow, rattle (instrument), bowed string instrument, sawing
+fold 3 rattle, slam, whoosh swoosh swish, hammer, fart, harp, coin (dropping), printer, boom, giggle
+fold 4 clapping, crushing, livestock farm animals working animals, scissors, writing, wind, crackle, tearing, piano, microwave oven
+fold 5 trumpet, wind instrument woodwind instrument, child speech kid speaking, drill, thump thud, drawer open close, male speech man speaking, gunshot gunfire, burping eructation, splash splatter
+fold 6 female singing, wind chime, dishes pots pans, scratching (performance technique), crying sobbing, waves surf, screaming, bark, camera, organ
 FSC22 train 0, 1, 2, 3, 4, 10, 11, 14, 16, 19, 20, 24, 25
@@
 test 5, 7, 15, 17, 21, 23, 26
-TABLE III: Classes in each partition.
+UrbanSound8K train 0, 1, 2, 4, 5, 7, 8
+val 3, 6, 9
+TAU 2019 train 2, 3, 4, 5, 7, 8, 9
+val 0, 1, 6
+GTZAN train 0, 1, 2, 6, 7, 8, 9
+val 3, 4, 5
+TABLE I: Classes in each partition.
 For ESC-50, random partitions are used from previous work for consistency [19].
-FSC22 has not been benchmarked for zero-shot learning before, so a new partition strategy is created with a random number generator.
-A modified YAMNet model [19] is trained using only the seen classes in FSC22.
-A learning rate of $10^{-3}$ is used and early stopping occurred at epoch 259 with an accuracy of $0.6833$.
-This model is used to retrieve the audio embeddings.
-The word embedding model is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/.
+All other datasets have not been benchmarked for zero-shot learning before, so new partitions are created using a random number generator.
+IV-B Embeddings
+A modified YAMNet model [20] - a YAMNet architecture with layers removed and an embedding layer added in before the output layer - is trained using only the seen classes in each dataset partition.
+For all datasets, the embedding layer size is 128, the learning rate is $10^{-3}$, and the batch size is 16.
+The models are trained with early stopping and a patience of 200.
+This model is then used to retrieve audio embeddings to train each zero-shot learning method.
+The word embedding model used is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/.
 Each method uses the same embedding vectors for each respective class and data sample.
-IV-A Bilinear Compatibility
-The bilinear compatibility network has been used to great success in environmental audio [18, 68, 19].
+IV-C Non-linear ALE
+The non-linear ALE compatibility network has been used to great success in environmental audio [48, 20].
 It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50.
 Therefore, it is important to include it in this comparison.
-The bilinear compatibility method has few hyperparameters and the parameters used by Xie and Virtanen [18] for ESC-50 were repeated in our work. FSC22 was tuned to train for 30 epochs.
-IV-B CADA-VAE
-Cross and distribution aligned variational autoencoder (CADA-VAE) [14] is a successful zero-shot learning method in computer vision, and for this reason it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
+The non-linear compatibility method has few hyperparameters and the parameters used by Xie et al.
+[48] for ESC-50 were repeated in our work for all datasets, except FSC22 which was tuned to train for 30 epochs.
+IV-D CADA-VAE
+Cross and distribution aligned variational autoencoder (CADA-VAE) [15] is a successful zero-shot learning method in computer vision.
+For this reason, it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
 Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets.
@@
 Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy.
-It was found that the large amount of hyperparameters made this method difficult and time consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
+It was found that the large amount of hyperparameters made this method difficult and time-consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
+The final training configuration was determined iteratively, guided by validation performance.
 The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error.
 The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22.
-The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs..
-The classifier has a learning rate of $10^{-3}$, a batch size of 32 and is trained for 23 epochs for ESC-50 and 20 epochs for FSC22.
-The loss of the VAE is mean absolute error and the loss of the classifier is negative log likelihood loss.
+The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs.
+The classifier uses a learning rate of $10^{-3}$ and a batch size of 32.
+It is trained for 23 epochs on ESC-50; 20 epochs on FSC22, UrbanSound8K and GTZAN; 30 epochs on TAU 2019; and 50 epochs on ARCA23K-FSD, with the number of epochs selected based on the size of each dataset.
+The loss of the VAE is a mean absolute error and the loss of the classifier is a negative log-likelihood loss.
 Both use an Adam optimiser.
 The size of the latent space is 64.
-IV-C LisGAN
+IV-E LisGAN
 Leveraging invariant side generative adversarial network (LisGAN) [16] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning.
-While there are some methodological concerns, these were modified to adhere to the zero-shot learning problem formulation and the core idea of the method remained.
-A conditional Wasserstein GAN is the main GAN method found in zero-shot learning.
+LisGAN is a transductive method while we consider only the inductive case in these experiments.
+The transductive component is removed and the core idea of the method remains.
+A conditional Wasserstein GAN [38] is the main GAN method found in zero-shot learning.
 However, despite efforts to improve the method, it is still notoriously difficult to train [57, 43].
 Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs.
-LisGAN showed to be significantly more stable and competitive amongst other zero-shot learning methods.
+LisGAN has shown to be significantly more stable and competitive amongst other zero-shot learning methods, even with the transductive component removed.
 Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune.
-These mostly were modified to suit the smaller ESC-50 and FSC22 datasets, as the datasets in computer vision are significantly larger.
-Both datasets used the same hyperparameters.
+These were modified to suit the smaller audio datasets, as the datasets in computer vision are significantly larger.
+All datasets used the same hyperparameters.
 LisGAN is used without retraining on unseen classes.
@@
 The generator, discriminator and classifier use Adam optimisers.
-The classifier uses a negative log likelihood loss.
-ESC-50 FSC22
-Val Fold 0 Val Fold 1 Val Fold 2 Val Fold 3 Test Val Test
-Bilinear $0.2460\pm 0.0074$ $0.5003\pm 0.0074$ $0.3345\pm 0.0105$ $0.2868\pm 0.0076$ $0.2257\pm 0.0105$ $0.3526\pm 0.0069$ $0.3343\pm 0.0169$
-LisGAN $0.2605\pm 0.0227$ $0.4039\pm 0.0220$ $0.2582\pm 0.0187$ $0.2150\pm 0.0195$ $0.2448\pm 0.0165$ $0.299\pm 0.0219$ $0.2142\pm 0.0416$
-CADA-VAE $\textbf{0.2915}\pm 0.0180$ $0.3980\pm 0.0181$ $0.2500\pm 0.0251$ $0.3013\pm 0.0251$ $0.2095\pm 0.0337$ $0.3179\pm 0.0137$ $0.3290\pm 0.0339$
-ZeroDiffusion $0.2846\pm 0.0111$ $\textbf{0.5279}\pm 0.0283$ $\textbf{0.3593}\pm 0.0189$ $\textbf{0.3342}\pm 0.0224$ $\textbf{0.3091}\pm 0.0278$ $\textbf{0.3975}\pm 0.0132$ $\textbf{0.3467}\pm 0.0200$
-TABLE IV: Average accuracy and standard deviation over ten randomly seeded runs for the traditional bilinear compatibility method and the three generative methods for ESC-50 and FSC22.
-Our method, ZeroDiffusion, outperforms the other methods in accuracy across the partitions of both datasets.
-The bilinear compatibility method has less variance than the three generative-based methods.
-IV-D ZeroDiffusion
-ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in this paper.
+The classifier uses a negative log-likelihood loss.
+ALE LisGAN CADA-VAE ZeroDiffusion
+ESC-50 fold 0 $0.2278\pm 0.0091$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2605}}\pm 0.0227$ $\textbf{0.2915}\pm 0.0180$ $0.2598\pm 0.0140$
+fold 1 $\textbf{0.4932}\pm 0.0093$ $0.4039\pm 0.0220$ $0.3980\pm 0.0181$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4665}}\pm 0.0194$
+fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2937}}\pm 0.0129$ $0.2582\pm 0.0187$ $0.2500\pm 0.0251$ $\textbf{0.3110}\pm 0.0082$
+fold 3 $0.2680\pm 0.0117$ $0.2150\pm 0.0195$ $\textbf{0.3013}\pm 0.0251$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2975}}\pm 0.0127$
+test $0.2340\pm 0.0167$ $\textbf{0.2448}\pm 0.0165$ $0.2095\pm 0.0337$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2343}}\pm 0.0118$
+average $0.3033\pm 0.0119$ $0.2765\pm 0.0199$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3102}}\pm 0.0240$ $\textbf{0.3138}\pm 0.0132$
+ARCA23K-FSD fold 0 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2420}}\pm 0.0032$ $0.2183\pm 0.0128$ $0.2398\pm 0.0125$ $\textbf{0.2879}\pm 0.0153$
+fold 1 $0.2876\pm 0.0028$ $0.3059\pm 0.0107$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3142}}\pm 0.0149$ $\textbf{0.3418}\pm 0.0150$
+fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3723}}\pm 0.0061$ $0.3086\pm 0.0121$ $0.3164\pm 0.0120$ $\textbf{0.4819}\pm 0.0188$
+fold 3 $\textbf{0.2546}\pm 0.0029$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2113}}\pm 0.0148$ $0.2036\pm 0.0101$ $0.1885\pm 0.0200$
+fold 4 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2283}}\pm 0.0020$ $0.2173\pm 0.0113$ $0.2126\pm 0.0140$ $\textbf{0.2833}\pm 0.0279$
+fold 5 $0.3205\pm 0.0050$ $0.3196\pm 0.0142$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3567}}\pm 0.0144$ $\textbf{0.4335}\pm 0.0321$
+test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2129}}\pm 0.0051$ $0.1772\pm 0.0068$ $\textbf{0.2414}\pm 0.0145$ $0.2081\pm 0.0261$
+FSC22 val $0.2650\pm 0.0088$ $0.2990\pm 0.0219$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3179}}\pm 0.0137$ $\textbf{0.3269}\pm 0.0113$
+test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2968}}\pm 0.0070$ $0.2142\pm 0.0416$ $\textbf{0.3290}\pm 0.0339$ $0.2775\pm 0.0071$
+UrbanSound8k $\textbf{0.4997}\pm 0.0071$ $0.3848\pm 0.0238$ $0.4741\pm 0.0808$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4961}}\pm 0.0777$
+TAU2019 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4377}}\pm 0.004$ $0.3878\pm 0.0188$ $0.3158\pm 0.0248$ $\textbf{0.4857}\pm 0.0395$
+GTZAN $0.5705\pm 0.0109$ $\textbf{0.5824}\pm 0.0287$ $0.5650\pm 0.0463$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.5803}}\pm 0.0281$
+Average $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3227}}\pm 0.0076$ $0.2936\pm 0.0187$ $0.3137\pm 0.0242$ $\textbf{0.3486}\pm 0.0221$
+TABLE II: Average accuracy and standard deviation over ten randomly seeded runs for the traditional non-linear ALE method and the three generative methods for the six chosen datasets.
+Our method, ZeroDiffusion, is similar or superior in almost all cases, with the exception of ESC-50 where no one method consistently performs the best.
+When averaging ESC-50 results, ZeroDiffusion is ahead of the other methods.
+The best accuracy in each row is bolded and the second-best accuracy is bolded in grey.
+The average value across all datasets is displayed in the final row, where ZeroDiffusion has the highest average accuracy and ALE reports the lowest overall variance.
+IV-F ZeroDiffusion
+ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in Section III-C.
 These experiments aim to determine the performance of this new method in the context of environmental audio.
-The existing success of diffusion in other areas of machine learning research motivate the creation and experimentation of this method.
-The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 2000 epochs.
-Fifty samples are generated from the diffusion model per unseen class.
-The bilinear compatibility network is trained with a learning rate of $10^{-4}$, batch size of 52, and 20 epochs.
-These hyperparameters were chosen using manual tuning.
+The existing success of diffusion in other areas of machine learning research motivates the creation and experimentation of this method.
+The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 50 epochs.
+The number of samples generated for each unseen class corresponds to the average number of samples per class in the dataset.
+For TAU 2019, this is 1,440 samples per class; UrbanSound8k is unbalanced but averages 873 samples per class; GTZAN is 100 samples per class; ESC-50 uses 40 samples per class; and finally, ARCA23K-FSD is unbalanced and averages 339 per class.
+We train the non-linear compatibility network with a learning rate of $10^{-4}$, batch size of 64, and 10 epochs.
+For the diffusion loss terms, reconstruction loss was weighted $1.0$, maximum mean discrepancy at $1.0$, variance loss as $0.1$, mean reconstruction loss at $0.2$, and cosine similarity at $2.0$.
+These hyperparameters were chosen using manual tuning across all validation sets.
+The same parameters were used for all experiments except for the synthetic dataset size, which changed relative to the dataset size.
 V Results
-The results from experiments with the four methods - bilinear compatibility, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50 and FSC22 are presented in Table IV.
-ZeroDiffusion increases the accuracy on the ESC-50 and FSC22 test partition by $36.95\%$ and $3.71\%$, respectively, from the baseline bilinear method.
-With the exception of the first validation fold in ESC-50, ZeroDiffusion has a higher accuracy than the other three zero-shot learning methods.
-This shows that ZeroDiffusion is a new state-of-the-art method for both ESC-50 and FSC22.
-Notably, ZeroDiffusion incorporates the bilinear model in its methodology.
-The difference between the methods is that ZeroDiffusion trains a diffusion model to generate synthetic unseen data and adds these samples to the training of the bilinear method.
-Because the diffusion model is the only addition, it is reasonable to conclude that the diffusion model is responsible for the improved accuracy.
-CADA-VAE has positive performance but is inconsistent across partition strategies.
-LisGAN has similar inconsistency issues but overall appears to have worse performance across partitions and datasets.
-This inconsistency continues when considering the standard deviation for each experiment.
-The generative methods have high variance compared to the baseline method.
-In the case of ZeroDiffusion, the higher average accuracy and variance are tradeoffs to consider.
-The accuracy improvement of ZeroDiffusion supports the use of diffusion models in the zero-shot learning field.
+Table II presents the results from experiments with the four methods - non-linear ALE, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k, TAU 2019 and GTZAN.
+ZeroDiffusion maintains an equally top result for UrbanSound8k and GTZAN, shared with ALE and LisGAN, while it outperforms other methods on TAU 2019.
+Notably, despite not outperforming all methods across all three smaller datasets, it is the only method to maintain a top result across all three.
+For the FSC22 dataset, the results are worse for ZeroDiffusion but it outperforms other methods on the validation set and underperforms on the test set.
+This indicates an issue with reliability for the new method; however, it is clear that the other presented methods also lack reliability.
+Overall, the accuracy improvement of ZeroDiffusion across most datasets and partitions supports the use of diffusion models in the field of zero-shot learning.
 Adjacent works with Stable Diffusion [44, 45] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning.
-Confusion matrices give more insight into model performance through the analysis of class predictions.
-The confusion matrices for one run of ESC-50 fold 0 and 1 are shown in Tables V and VI, respectively.
-In fold 0 of the ESC-50 dataset, the helicopter and mouse click classes have the highest accuracy at 80% and 62.5%, respectively.
-The helicopter class is predicted at least 70% of the time for four different classes, resulting in a hub around the helicopter class.
-A hub is a class where predictions across classes are clustered, i.e.
-the model is much more likely to predict the hubs than other classes.
-In zero-shot learning, the hubness problem has been identified as a major issue preventing classes from being predicted, resulting in a low accuracy.
-In fold 0, smaller hubs exist around drinking, brushing teeth and mouse clicking.
-Other classes are not predicted often.
-The difference between the hub classes and other classes does not appear to have a clear intuitive explanation.
-This information could lead to avenues for improvement.
-Fold 1 of ESC-50 has a significantly higher accuracy than fold 0.
-The highest class accuracy is siren at 80%, however all but the clapping class have at least 25% accuracy.
-Clapping is predicted a total of three times across all clips in this fold, and all three are misclassified hand saw clips.
-Other classes are predicted more consistently, indicating that the hubness problem isn’t as prevalent in the training of this partition of the dataset.
-The confusion matrix in fold 1 is sparse, where classes are only misclassified as a small number of other classes.
-This phenomenon is also found in traditional supervised learning, where the target class is misclassified as classes with similar features.
-In zero-shot learning, it is the machine learning-based semantic embeddings that are being compared, and so intuition from traditional supervised learning may not apply.
-Some misclassifications appear intuitive, such as the case where the sound of a washing machine is incorrectly labeled as a helicopter.
-Both classes are loud and mechanical, motivating a potential misunderstanding from the model.
-Fireworks are misclassified as mouse clicking in fold 0, and the reason may not initially be intuitive, but both sounds are comprised of short, sharp noises.
-There is a difference in loudness, but the audio data is normalised before input to the network and so this difference may be reduced.
-While there are some intuitive misclassifications, there are also non-intuitive misclassifications.
-Pig is incorrectly labeled as brushing teeth, and door knocking is mislabeled as rain.
-To gain a deeper understanding, dimensionality reduction with t-distributed stochastic neighbour embedding (t-SNE) [72] on the audio embeddings produces Figure 3 for fold 0 and Figure 4 for fold 1.
-In fold 0, pig (class 2, dark purple) is spread across a wide area, including overlapping with brushing teeth (class 27, dark blue).
-This relationship gives insight into the non-intuitive misclassification of pig.
-Fold 1 has a significantly higher accuracy than fold 0, as shown in Table IV.
-The t-SNE graph for fold 1 shows more distinct clusters than fold 0, indicating that this clustering property of the audio embeddings is important in correctly classifying unseen classes.
-This is reinforced when considering the clapping class, which both is the poorest performing class in fold 1 and has the greatest spread across the t-SNE graph.
-The t-SNE graphs reveal insights into the hubness problem, and asymmetry in the incorrect predictions further highlights the problem.
-Pig is mislabeled as brushing teeth more than 50% of the time, while brushing teeth is never mislabeled as pig.
-This shows that certain classes are more likely to be predicted than others.
-The classes in fold 0 that are frequently predicted as the helicopter class are closer together in the t-SNE graph, showing their similarity in the embedding space.
-This similarity is a potential cause of the hubness problem.
-Class index 38 29 35 3 40 2 27 46 31 48
-38 1 9 0 3 0 3 9 0 14 1
-29 1 18 0 0 1 0 18 0 1 1
-35 1 0 6 0 30 1 1 1 0 0
-3 0 0 5 0 29 5 1 0 0 0
-40 0 0 6 0 32 0 0 1 0 1
-2 0 6 3 0 3 5 21 0 0 2
-27 2 34 0 1 0 0 1 0 2 0
-46 0 0 2 0 28 2 0 8 0 0
-31 1 5 0 0 2 0 6 0 25 1
-48 4 6 0 0 2 0 8 1 15 4
-TABLE V: Confusion Matrix for fold 0 of the ESC-50 dataset.
-Rows represent the true class, and columns represent the predictions.
-Many predictions are clustered around class 40, with smaller hubs on classes 29, 27 and 31.
-Figure 3: ESC-50 fold 0 audio embeddings with t-SNE dimensionality reduction.
-Class 40 (helicopter), 46 (church bells), 35 (washing machine) and 3 (cow) are all concentrated to the right side of the scatterplot.
-The confusion matrix in Table V shows that all of these classes are classified as helicopter a majority of the time.
-Class index 39 36 42 13 32 22 19 49 26 21
-39 10 0 0 0 9 0 2 3 0 16
-36 0 24 3 4 0 0 6 1 0 2
-42 0 0 32 3 0 0 0 2 3 0
-13 0 4 7 14 1 0 5 0 1 8
-32 2 0 0 0 23 0 2 13 0 0
-22 14 1 0 1 3 0 14 6 1 0
-19 0 5 10 0 0 0 22 0 0 3
-49 1 3 0 1 0 3 0 11 8 13
-26 4 0 2 4 0 0 0 4 15 11
-21 2 0 0 1 1 0 0 0 8 28
-TABLE VI: Confusion matrix for fold 1 of the ESC-50 dataset.
-Rows represent the true class, and columns represent the predictions.
-There are small hubs on classes 42, 19, 26 and 21 and all classes are at least partially correctly identified, except for class 22 which is never correctly classified.
-Figure 4: ESC-50 fold 1 audio embeddings with t-SNE dimensionality reduction.
-Fold 1 exhibits more defined class clusters than fold 0.
-Class 22 (clapping, dark blue) has a greater spread, and is the poorest classified class as shown in Table VI.
-t-SNE graphs give a deeper insight into issues surrounding zero-shot learning, and can indicate avenues for improvement.
-The relationship between the clustering properties of the audio embeddings and model accuracy show that embeddings have an effect on the performance of a zero-shot learning model.
-More defined clustering in the audio embeddings produces higher accuracy, and hubs in the confusion matrix are reflected in the t-SNE diagram.
-The confusion matrices highlight the abilities of the model beyond an accuracy value.
-The sparsity of the matrices demonstrate that the model recognises the similarity of clips from the same class, and when it correctly maps these to the correct class the performance is very high.
+ZeroDiffusion is agnostic to the specific data type and works entirely with embeddings, similarly to ALE, and therefore is a more flexible method than the image domain-specific Stable Diffusion methods.
+V-A Comparison of Dataset Sizes
+ARCA23K-FSD is a significantly larger dataset than any of the other datasets, and as such allows for a deeper 7-fold analysis.
+ZeroDiffusion outperforms the other methods in almost every case, and in the cases of folds 0, 2, 4 and 5 it has a significant improvement over any other method.
+This result indicates that ZeroDiffusion may draw benefits from the larger dataset, with richer embeddings produced by the diffusion model as a result of more classes available.
+However, despite the overall success, fold 3 extremely underperforms and could provide an avenue for improvement.
+Whilst ESC-50 is not as large as ARCA23K-FSD, its large class size allows for 5-fold partitioning.
+However, the results for this dataset are extremely unstable when considering the best-performing methods.
+No one method consistently outperforms this dataset, and all methods outperform the others on at least one partition.
+To gain a better understanding, the average across all folds is calculated and ZeroDiffusion marginally has the highest accuracy.
+An indication of why this is happening may be revealed in the average number of audio clips per class in each dataset.
+ESC-50 has the lowest with 40 samples per class.
+This may directly result in the method failing to generalise.
+FSC22 is the next largest dataset per class with 75 samples, and this dataset similarly experienced issues with performance.
+Of the smaller datasets, TAU 2019 has 1,440 per class, the highest of any dataset in the experiments, and ZeroDiffusion produced a significant improvement over other methods.
+ARCA23K-FSD has on average 339 samples per class, a reasonable increase over both ESC-50 and FSC22, and reports good results.
+Given this is an unbalanced dataset, this may provide insight into why ZeroDiffusion underperforms on some folds in ARCA23K-FSD.
+V-B Ablation
+Notably, ZeroDiffusion incorporates the non-linear model in its methodology.
+The ALE results therefore represent an ablation of ZeroDiffusion without the diffusion-generated data component.
+The only other changes present are the weight decay term on the classifier ($10^{-4}$ for ALE and $10^{-5}$ for ZeroDiffusion) and the loss function (WARP for ALE and cross-entropy for ZeroDiffusion’s classifier); however, tests were done to align these with poorer results on both ALE and ZeroDiffusion across datasets when changing either or both terms.
+Given all of this, is it reasonable to conclude that the addition of a diffusion model to generate synthetic data for classifier training is responsible for the improved accuracy between ALE and ZeroDiffusion.
+V-C Variance
+Standard deviation is recorded for each experiment, and it is clear that across runs the generative-based methods have a high standard deviation.
+Meanwhile, ALE has a relatively low standard deviation.
+This indicates that generative-based methods, including ZeroDiffusion, are more variable and potentially unreliable compared to the more stable ALE method.
+Additionally, generative methods typically have a wide range of hyperparameters.
+Methods such as ALE have a low number of hyperparameters, making them user-friendly and simple to apply to new problems.
 VI Discussion
+Given the results of this four-model and six-dataset study, three areas are considered: model consistency and the importance of consistency in zero-shot learning; the impact of dataset size; and the issue of theoretical guarantees in zero-shot learning.
+VI-A Consistency of Methods
 In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples.
-Therefore, consistency across training runs is important.
-While ZeroDiffusion has a superior accuracy-based performance, the bilinear method has superior consistency.
-This is an important consideration when choosing an appropriate method for zero-shot learning in practical applications.
-The hubness problem is an issue when considering practical usage because of the poor accuracy on specific individual classes that are overshadowed by the hubs.
-Many of these specific classes are predicted at a rate worse than a random model.
-Hubness appears to correlate to overall model performance, as seen when comparing the confusion matrices and overall accuracy of two partitions of the ESC-50 dataset.
-One key avenue in reducing the hubness problem is first understanding why the different partitions used in this work had varying levels of hubness.
-Partitions can be crafted for this purpose to identify the effect of different criteria on hubness.
-One question is whether the unseen classes, seen classes, or both, influence the performance.
-If particular characteristics in the seen class set influence the hubness problem, then the seen class set can be supplemented with classes exhibiting these characteristics.
-A more likely scenario is that both influence the performance, and the seen classes used in training may need to be curated for the characteristics of the specific unseen set in the given problem.
-Zero-shot learning lacks theoretical foundations, with methods developing through empirical results.
-There is little understanding of upper bounds to accuracy for a given unseen and seen class sets.
+Therefore, consistency across training runs and datasets is important for building trust in the model.
+The issue of the black-box nature of machine learning models hinders application to key global problems, and the added uncertainty of model performance further exasperates these problems.
+When considering the standard deviation of each result in Table II, there is a high variance trend amongst the generative methods (LisGAN, CADA-VAE, ZeroDiffusion), with CADA-VAE reaching a high 8.08% standard deviation on the UrbanSound8k dataset and ZeroDiffusion not far behind with a 7.77% deviation.
+Meanwhile, the highest standard deviation for ALE is around 1% on ESC-50 partitions and GTZAN.
+This is a serious problem for some domains and indicates a need to develop more stable generative methods to make use of the improved accuracy of these methods.
+Between datasets, there is also an inconsistency, as noted with ESC-50 having top results across partitions from every method used in the presented experiments.
+ZeroDiffusion has high accuracy on ARCA23K-FSD; however, it suffers issues with fold 3 and the test fold despite large improvements on the other folds.
+Prediction of model performance across partitions and datasets can better inform the application and improvement of these methods.
+VI-A1 Data Diversity
+From the results presented, larger datasets typically faired better with ZeroDiffusion.
+TAU 2019 may have only ten classes, but with 1,440 samples per class it represents a large audio dataset and this appears to result in an accuracy improvement over other methods.
+The smaller two datasets, ESC-50 (40 samples per class) and FSC22 (75 samples per class) experienced instability in performance across methods.
+ARCA23K-FSD provided a balance between the number of classes (70) and the number of samples per class (on average 339 samples), and the methods were more consistent in their performance, with ZeroDiffusion dominating.
+The dataset size, both in the number of classes and samples per class, may influence which zero-shot classification method is more suitable for individual scenarios and tasks.
+This is an important consideration when applying zero-shot learning in practice.
+VI-A2 Theoretical Guarantees
+Zero-shot learning lacks theoretical foundations, with methods developing through empirical results, as done in the presented work.
+There is little understanding of upper bounds on accuracy for a given set of unseen and seen classes.
 A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes.
-A trivial example is if the seen class set is the empty set.
-A deeper understanding of these requirements may be an avenue to reducing hubness, and therefore improving overall classification performance.
+A trivial example is if the seen class set is the empty set, then there is no information to learn.
+One indication of these theoretical guarantees is the differences between dataset sizes in the presented experiments and the effect on model stability.
+A deeper understanding of these requirements may improve overall classification performance.
 VII Conclusion
-A comparison of three generative methods and the baseline method in environmental audio zero-shot learning has been presented.
-Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly or worse than the baseline, and present higher variance.
-The third generative method, ZeroDiffusion, is a novel diffusion-based method for zero-shot learning.
-It shows a higher accuracy performance than the other three methods, thus producing a new state-of-the-art result in the domain of environmental audio on ESC-50 and FSC22.
-Diffusion shows potential in zero-shot learning to improve the performance across domains.
-Future work could include benchmarks for diffusion-based zero-shot learning models on datasets in areas other than environmental audio.
-An analysis of confusion matrices and t-SNE graphs of the audio embeddings reveals the impact of audio embedding similarities on the hubness problem and model performance.
-This analysis is valuable since these similarities may not be intuitive from the raw audio data.
+This study presents a comparison of three generative methods and a baseline method in zero-shot environmental sound classification.
+Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly to the baseline and present higher variance.
+The third generative method, ZeroDiffusion, is our novel diffusion-based method for zero-shot learning.
+It has a higher accuracy performance than the other three methods in most cases on the six chosen datasets, with a new state-of-the-art result on ARCA23K-FSD and TAU Urban Acoustics 2019, and reasonable results on ESC-50, FSC22, UrbanSound8k and GTZAN.
+The inclusion of TAU 2019 and GTZAN shows that the method can perform in applications other than environmental sound classification, with acoustic scene classification and music genre classification represented by these datasets.
+Future work on diffusion-based zero-shot models would benefit from considering other domains, such as image classification.
 Acknowledgements